### v1

# Look Twice Before You Answer: Memory-Space Visual Retracing for Hallucination Mitigation in Multimodal Large Language Models

Abstract
Despite their impressive capabilities, Multimodal Large Language Models (MLLMs) are susceptible to hallucinations, especially assertively fabricating content not present in the visual inputs. To address the aforementioned challenge, we follow a common cognitive process - when one’s initial memory of critical on-sight details fades, it is intuitive to look at them a second time to seek a factual and accurate answer. Therefore, we introduce Memory-space Visual Retracing (MemVR), a novel hallucination mitigation paradigm that without the need for external knowledge retrieval or additional fine-tuning. In particular, we treat visual prompts as supplementary evidence to be reinjected into MLLMs via Feed Forward Network (FFN) as “key-value memory”, when the model is uncertain or even amnesic about question-relevant visual memories. Comprehensive experimental evaluations demonstrate that MemVR significantly mitigates hallucination issues across various MLLMs and excels in general benchmarks without incurring added time overhead, thus emphasizing its potential for widespread applicability.  [^1] Code can be available at https://github.com/1zhou-Wang/MemVR.
614
Figure 1: MemVR demonstrates strong performance across seven benchmarks spanning various domains (Left), with particularly outstanding results on the MME benchmark (Center). Additionally, MemVR is compared with contrastive decoding schemes, standing out for its ability to alleviate hallucinations using just a single inference, making it a more eco-friendly solution (Right).
1 Introduction
Multimodal Large Language Models (MLLMs), due to their formidable capacity to comprehend visual inputs, have emerged as indispensable tools in computer vision (Koh et al., 2024) and natural language processing (Tu et al., 2023) to tackle numerous visual tasks and facilitate complex visual question answering (VQA). Nevertheless, MLLMs still exhibit certain limitations, i.e., the so-called “hallucination” phenomenon (Huang et al., 2024b; Zheng et al., 2024). Specifically, MLLMs frequently generate descriptions inconsistent with user-provided visual inputs, such as incorrectly outputting non-existent things in the image or conflicting judgments. This flaw poses significant risks to the reliability of MLLMs as trustworthy assistants (Yu et al., 2024b), particularly in safety-critical applications (Zou et al., 2023) (e.g., clinical healthcare (Lin et al., 2024) and autonomous driving scenarios (Ding et al., 2024)). While the exact reasons for hallucinations in MLLMs are not fully understood, one possible factor could be the imbalance between their understanding of visual and textual information. Typically, MLLMs encode images into vision tokens via CLIP (Radford et al., 2021), which are fed along with text tokens into LLMs for decoding. LLMs excel at text comprehension but struggle with visual information perception and memory, where differences in information density between modalities may cause inconsistencies during decoding.
Numerous methods have been proposed to mitigate hallucinations in MLLMs, and general studies can be broadly categorized into three streams: (i) Retrieval-Augmented Generation (RAG) (Shuster et al., 2021; Caffagni et al., 2024) that incorporates knowledge from external databases to mitigate the problem of “hallucination”, as well as (ii) through extra Fine-tuning (Yu et al., 2024b) to enhance the self-consistency of generation, and (iii) Contrastive Decoding (CD) strategy, which do not involve extra training. Specifically, RAG and fine-tuning patterns typically employ external knowledge retrieval or robust instruction-tuning datasets to post-hoc debias (Yang et al., 2024; Liu et al., 2023a), which inevitably introduces substantial computational overhead or storage requirements. For example, some approaches(Yin et al., 2023; Yu et al., 2024a) have fine-tuned models using high-quality visual instructions generated by advanced automated annotation tools including GPT-4 (OpenAI, 2023).
CD-based methods (Li et al., 2023b; Shi et al., 2024) represent a simpler and more efficient way to mitigate hallucinations than RAG and fine-tuning based methods. Particularly, CD-based hallucination mitigation methods usually modulate the logits of the next token prediction through contrast manner or penalty mechanisms. As illustrated in Figure 2 (a), DoLa (Chuang et al., 2023) enriches factual knowledge via layer-wise contrasting and reduces the generation of incorrect facts in LLMs. In MLLMs, VCD (Leng et al., 2024) amplifies the language priors by adding Gaussian noise to the visual inputs, thereby reducing over-reliance on statistical biases and single-modal priors through contrasting output distributions from original and distorted visual inputs as in Figure 2 (b). This perturbation of original inputs requires task-specific design, inevitably doubling inference costs. More critically, contrastive distributions are agnostic to visual and instructional nuances, which may not always amplify the intended hallucinations, occasionally introducing potential noise into CD.
Figure 2: The conventional paradigm for hallucination mitigation, employing contrastive decoding methods such as DoLa (a) and VCD (b), is compared to our proposed MemVR strategy (c).
Method 20-Token Len 50-Token Len 80-Token Len
LLaVA-1.5 1880.3  $\downarrow$×$\times$1.0 3617.6  $\downarrow$×$\times$1.0 5256.6  $\downarrow$×$\times$1.0
+ VCD 4537.4  $\uparrow$×$\times$2.4 7690.8 $\uparrow$×$\times$2.1 11569.3 $\uparrow$×$\times$2.2
+ OPERA 6242.7  $\uparrow$×$\times$3.3 12672.3  $\uparrow$×$\times$3.5 19247.2  $\uparrow$×$\times$3.7
+ MemVR 1861.7  $\uparrow$×$\times$1.0 4000.9  $\uparrow$×$\times$1.1 5545.5  $\uparrow$×$\times$1.1
Table 1: Efficiency Comparisons for generating different length tokens, using an NVIDIA-A40 GPU. Inference time (ms) of different methods is recorded.
In this work, we delve into the challenges of hallucination mitigation in MLLMs and address the shortcomings of CD-based approaches. Our research is grounded in a common cognitive process: when the initial memory of certain critical visual details fades, it is intuitive to look at them for the second time to search for the accurate answer (O’regan, 1992; Ballard et al., 1995; Horowitz & Wolfe, 1998). Different from visual contrastive decoding strategies that alleviate hallucinations by diminishing language priors (Leng et al., 2024; Qu et al., 2024; Park et al., 2024), we propose a novel Memory-space Visual Retracing (MemVR) method that mitigates hallucinations through supplementing visual evidence, akin to the two sides of a coin. MemVR, as shown in Figure 2 (c), is an architecture-agnostic, plug-and-play solution that re-injects visual features into an intermediate layer suffering from vision-related memory lapse with only one regular inference. This novel hallucination mitigation paradigm in terms of efficiency, and its inference cost and performance are optimal compared to previous studies as listed Table 1. It’s a game-changer for effectiveness and efficiency. Through extensive experiments on multimodal hallucination benchmarks, as well as GPT-4o [^2] GPT-4o-2024-08-06: https://platform.openai.com/docs/models/gpt-4o. evaluations, we show the comprehensive performance improvements of MemVR in hallucination mitigation and general capabilities. Our contributions can be summarized as follows:
  • We propose MemVR, a novel training-free hallucination mitigation paradigm that effectively alleviates hallucinations in MLLMs. In contrast to previous methods, which primarily focus on eliminating biases of language priors, MemVR seeks to replenish question-relevant visual clues towards more evidential responses, which signifies the other side of the coin.
  • We design a dynamic premature layer injection strategy with visual retracing in MLLMs, mimicking human intuitive thinking to revisit image features for self-consistency and credible answers when pivotal memories are scrambled. Furthermore, we theoretically demonstrate that visual retracing can effectively diminish hallucinations from an information-theoretic perspective.
  • Comprehensive experimental results demonstrate the effectiveness of MemVR in mitigating hallucinations and enhancing general cognitive and perceptual performance, as well as its high efficiency. Our research will make a substantial contribution to trustworthy multimodal intelligence.
2 Related Work
MLLMs and Challenges. In recent years, MLLMs have made remarkable progress, particularly as they have evolved from the foundations laid by Vision Language Models (VLMs). Early based on BERT-style language decoders (Devlin, 2018), which achieved initial cross-modal integration by combining visual and textual data (Li et al., 2022). Leveraging open-source Large Language Models (LLMs) such as LLaMA families (Touvron et al., 2023), MLLMs (Alayrac et al., 2022; Wu et al., 2024) have demonstrated enhanced adaptability across a range of visual language tasks, leading to a more profound ability to interpret the world. Models like LLaVA (Liu et al., 2024), Qwen-VL (Bai et al., 2023), and GLM4V (Wang et al., 2023) have further advanced this field, enabling users to interact with these agents using both image and text prompts. These models adhere to two critical training phases: pre-training feature alignment and instruction fine-tuning, ensuring they better comprehend the format of instruction inputs (Yin et al., 2024). However, despite their impressive performance in many areas, MLLMs still suffer from hallucination issues. Thus, in this paper, we primarily conducted experiments and analysis on these three representative models.
Hallucinations in MLLMs. Before LLMs, hallucination in NLP was mainly seen as generating nonsensical or deviant content. In MLLMs, “hallucination” is defined as the model generating content inconsistent with the provided image. This issue stems significantly from inadequate alignment among modalities. Several methods have been explored to mitigate hallucinations in MLLMs. Early efforts focused on fine-grained modality alignment (Rohrbach et al., 2018) and reducing co-occurrence biases (Kim et al., 2023) in small-scale VLMs, but these approaches struggle to scale with MLLMs. More recent strategies involve hallucination-targeted datasets for fine-tuning (Gunjal et al., 2024), post-hoc revisors (Zhou et al., 2024), and adopting RLHF (Yu et al., 2024b). While effective, these methods are resource-intensive. CD-based approaches (Chuang et al., 2023; Leng et al., 2024) adjust the decoding distribution to mitigate hallucinations, but it does not consistently improve performance. Compared with them, our MemVR stands as “a paradigm of effectiveness and efficiency” in hallucination mitigation, effortlessly enhancing performance without extra training.
3 Methodology
In this section, we first formulate the generation process of MLLMs to facilitate a clearer understanding of our MemVR. Moreover, we introduce our hypothesis about the causes of hallucinations and discuss visual retracing and its dynamic strategy. Further, we conduct theoretical analysis.
3.1 MLLMs Generation Process
Typically, MLLMs are composed of a visual encoder ${C}_{v}$, a text embedding layer, $L$ number of transformer layers, and an affine layer $\zeta(\cdot)$ which predicts the distribution of the next token, with diverse modalities as inputs, e.g., images and text. Regardless of specific architectural variations, MLLMs commonly employ ${C}_{v}$ to extract vision tokens from the raw image and project it into modality-shared feature space via an MLP or Q-Former module (Wadekar et al., 2024). Aligned vision tokens are represented as $X_{v}=\{x_{1},x_{2},\dots,x_{n_{v}}\}$ and used as part of the LLM input alongside the text tokens $X_{q}=\{x_{n_{v}+1},x_{n_{v}+2},\dots,x_{n_{v}+n_{q}-1}\}$ that are embedded from tokenized text input by the embedding layer. Subsequently, the vision and text tokens are concatenated as the final input sequence and we denote it as $\{x_{i}\}_{1}^{t_{n}-1}$ where $t_{n}=n_{v}+n_{q}$, which is then fed into successive transformer layers. We denote the output of the $l$-th layer as $h_{t}^{(l)}$. Then, the vocabulary head $\zeta(\cdot)$ is used to predict the probability of the next token $x_{t}$ among the vocabulary set $\mathcal{X}$ as follows,
$p(x_{t}\mid x_{<t})=\mathrm{softmax}(\zeta(h_{t}^{(L)}))_{x_{t}},x_{t}\in\mathcal{X}.$ (1)
3.2 What Causes Hallucinations
A hypothesis on the cause of hallucinations. Informed by the phenomenon of catastrophic forgetting (Zhai et al., 2023) in MLLMs, we argue that the capabilities of LLMs to comprehend and memorize different modalities are quite distinct. Taking image and text inputs as an example, since an image possesses a much higher information density than a piece of text, it is reasonable to assume that LLMs struggle to understand and memorize vision tokens compared to text tokens, prone to fantasies.
Uncertainty quantification. Following the DoLa (Chuang et al., 2023), we compute the probability of the next token via the vocabulary head $\zeta$ on each layer during reasoning. Then, we introduce an entropy-based metric (Farquhar et al., 2024) to quantify the output uncertainty as $u=\sum-p_{k}\log p_{k}/\log{K}$, where $\{p_{k}\}_{i=1}^{K}$. With uncertainty, we make an important assumption.
Assumption 3.1.
LUFH: The Lower the Uncertainty, the Fewer the Hallucinations to be generated.
We define $\gamma$ as the threshold that separates high uncertainty from low uncertainty in predictions. The candidate layer with uncertainty exceeding $\gamma$ is termed a premature layer. The proof of Assumption 3.1 (LUFH) is provided in Section 4. Based on the hypothesis, we aim to complement visual evidence in MLLMs to eliminate the hallucination caused by visual forgetting. In the following, we discuss our motivation, how MemVR is implemented, and why it can work in Section 3.3 and 3.4.
3.3 Relationship Between Hallucinations and Uncertainty
As findings of Chen et al. (2024) in LLMs: “incorrect tokens generally exhibit higher entropy than correct ones”, we also observe this phenomenon in MLLMs (visualization cases are shown in Appendix C.3). This implies the effectiveness of entropy-based metrics for detecting hallucinations. In this work, we use uncertainty as the metric. We present our in-depth findings in this section.
Figure 3: Uncertainty of different early layers to predict the next token. Rows denote indices of the early layers, and column names are decoded tokens in each step. Uncertainty distribution is dynamic.
Finding #1: In the context of tokens involving objects, attributes or relations, uncertainty is high.
We conduct preliminary analysis with 32-layer LLaVA-1.5-7B. Specifically, we compute the uncertainty in the output distributions $p^{(l)}(\cdot\mid x_{<t})$ of early exiting layers. Figure 3 shows the uncertainty scores of different early layers when decoding the answer, we can observe that the computed uncertainty remains relatively high in later layers when predicting key entity objects, attributes, or relations, such as wooden, filled, and food in Figure 3. This phenomenon suggests that LLM is still uncertain about its predictions in the last few layers and may inject more factual knowledge into the predictions. On the other hand, when predicting function words and those tokens copied from the question, e.g., image, a, with, we observe that the uncertainty becomes very low from the middle layers. This finding implies that the model is deterministic for easy-to-predict tokens at the intermediate layer and keeps the distribution of outputs almost constant at higher layers, however, it is more uncertain for difficult-to-predict key tokens and may constantly change its predictions until the final layer.
Finding #2: The uncertainty distribution of different tokens is dynamic during the generation.
As can be seen in Figure 3, when visual factual knowledge is required for the prediction of the next token, such as wodden, filled, food, most layers of LLaVA are of higher uncertainty, and appear to shift the prediction of the later layers. The prediction of the next tokens such as quantity, target category, color, relation, etc. requires evident visual knowledge, but not for function words or commas
From a qualitative perspective, intermediate values of uncertainty reflect the thinking or revision process of MLLMs, while too high uncertainty implies that the model is confused and relies on random guesses. For complex input images, the uncertainty of the output distributions tends to rise, which can propagate through subsequent layers, resulting in incorrect predictions, or gibberish.
Figure 4: The illustration of how dynamic premature layer injection works.
3.4 Visual Retracing and Its Dynamics
Motivated by the above findings, we propose re-injecting visual evidence during elevated uncertainty in the model’s reasoning. This strategy treats visual tokens as anchors to recalibrate off-target predictions and reduces uncertainties in object, attribute, relationship tokens. Experimental results also demonstrate that our method reduces uncertainty and alleviates hallucinations as shown in Figure 9. We term this schema of re-injecting visual evidence as “visual retracing” that is to be elaborated further in Section 3.4.1. Further, we design a dynamic injection strategy, detailed in Section 3.4.2, ensuring timely visual evidence when generating uncertain visual-reliant tokens (Figure 4).
3.4.1 FFN with Visual Retracing
We introduce the implementation of our proposed memory-space visual retracing method. Previous study (Geva et al., 2021) has found that FFN acts as a key-value memory storing factual knowledge. Inspired by the fact that FFN executes analogous retrieval from its key-value memory, we consider “visual retracing” to serve as a simplified and efficient information re-retrieval process (Jie et al., 2024). Concretely, given an hidden token $x\in\mathbb{R}^{d}$ and vision features $z_{v}=(z_{v,1},\dots,z_{v,n_{v}})^{\top}\in\mathbb{R}^{n_{v}\times d^{\prime}}$, $\mathcal{K}_{\propto}\left({z}_{v,i}\right),\mathcal{V}_{\propto}\left({z}_{v,i}\right)\in\mathbb{R}^{d}$ are the key and value corresponding to ${z}_{v,i}$, visual retracing follows,
${\operatorname{FFN}}_{\circlearrowright}^{(l)}({x}\propto z_{v})=(1-\alpha)\operatorname{FFN}^{(l)}(x)+\alpha\underline{\Lambda}_{\circlearrowright}^{(l)}({z}_{v}\mid{x}),$ (2)
where $\alpha$ denotes the injection ratio of visual memory (proportional to image complexity), FFN comprises two FC layers with non-linear activation in between and can be formulated as $\operatorname{FFN}({x})=\phi\left({x}{W}_{1}\right){W}_{2}^{\top}$, $\phi$ is activation function like ReLU or SiLU (Liu et al., 2020). Separately, the weight matrices can be rewritten as: ${W}_{1}=\left({k}_{1},{k}_{2},\ldots,{k}_{m}\right),{W}_{2}=\left({v}_{1},{v}_{2},\ldots,{v}_{m}\right)^{\top}$, in which ${k}_{i}\in\mathbb{R}^{d}$ and ${v}_{i}\in\mathbb{R}^{d}$ are entries of key and value, respectively. Thus, FFN can be interpreted as using input ${x}$ as the query to calculate its similarity with keys to search for matching values. Analogously, we consider a simple and efficient retrieval process for visual retracing on $l$-th premature layer as,
$\underline{\Lambda}_{\circlearrowright}^{(l)}({z}_{v}\mid{x})=\sum\nolimits_{i=1}^{n_{v}}\phi\left(\left\langle{x},\mathcal{K}_{\propto}\left({z}_{v,i}\right)\right\rangle\right)\cdot\mathcal{V}_{\propto}\left({z}_{v,i}\right).$ (3)
From the perspective of FFN, we regard the $\left(\mathcal{K}_{\propto}\left({z}_{v,i}\right),\mathcal{V}_{\propto}\left({z}_{v,i}\right)\right)$ as new key-value entries to supplement vision-related evidence. $\mathcal{K}_{\propto}$ and $\mathcal{V}_{\propto}$ are key functions for aligning the dimension between vision features $z_{v}$ and the hidden token $x$. In MLLMs, $\mathcal{K}_{\propto}$ and $\mathcal{V}_{\propto}$ are regarded as visual projectors, e.g., MLPs or Q-Former (Li et al., 2023a), for aligning features of different modalities.
3.4.2 Dynamic Premature Layer Injection
To magnify the effectiveness of visual retracing, the optimal premature layer should ideally be the layer most uncertain about probable answers to visual questions. In practice, we consider that the uncertainty of a candidate layer exceeding the threshold $\gamma$ warrants visual retracing. Inspired by the fact that early exit patterns (Teerapittayanon et al., 2016; Elbayad et al., 2020; Schuster et al., 2022) have proven effective in directly employing the language heads $\zeta$ to the hidden states of the middle layers, even without a special training process (Kao et al., 2020), we compute the uncertainty of the next token probability on the early layers for reasoning. As our Finding #2, we utilize layer-specific uncertainty to allow for dynamic premature layer injection at each time step as illustrated in Figure 3.
Algorithm 1 Dynamic Injection Strategy
1:$h_{t},{z}_{v}$ denote hidden states and visual evidence. We set trigger = True.
2:for $l$ = 1 to $L-1$ do
3:     $p^{(l)}=\mathrm{softmax}(\zeta(h_{t}^{(l)}))_{x_{t}}$
4:     $u^{(l)}={\sum-p^{(l)}\log p^{(l)}}/{\log{K}}.$
5:     if trigger==True and $u^{(l)}>\gamma$ then
6:         Execute $\underline{\Lambda}_{\circlearrowright}^{(l+1)}({z}_{v}\mid{h_{t}^{(l+1)}})$
7:         Select ${\operatorname{FFN}}_{\circlearrowright}^{(l+1)}({h_{t}^{(l+1)}}\propto z_{v})$
8:         trigger=False # only once
9:     else
10:         Select ${\operatorname{FFN}}^{(l+1)}({h_{t}^{(l+1)}})$
11:     end if
12:end for
For MLLMs with different numbers of layers, we partition the layers into several (typically three) buckets according to the total layers for sensible visual retracing triggers, as detailed in Appendix C.1. This dynamic injection strategy identifies desirable premature layers for visual retracing based on output uncertainty, thus better leveraging insights from different layers.
Static Fixed Layer Injection. In addition to the dynamic premature layer injection strategy, another more straightforward strategy worth considering is to perform a brute-force experiment on all possible early layers using a validation set and selecting the layer with the best average performance. We refer to this simple strategy as MemVR-static. However, the MemVR-static approach presents two limitations: (I) it requires more extensive hyperparameter tuning across layers, and (II) the optimal layer is highly sensitive to data distribution, necessitating an in-distribution validation set. In contrast, our proposed dynamic layer injection strategy mitigates these challenges by reducing the layer search space and improving robustness without depending on in-distribution validation. Empirical comparisons between our MemVR using dynamic and static strategies are provided in Section 4.2 and Table 4.
3.5 Theoretically Understanding Why MemVR Works
In order to gain further insight into the reasons behind the effectiveness of MemVR in mitigating hallucinations and its robust performance on general benchmarks, we attempt to explain these phenomena via the following three theorems from an information-theoretic perspective.
Theorem 3.1.
Let ${H}_{vq}$ be the hidden states of FFN and $\hat{H}_{vq}$ be after reinjection of visual evidence $Z_{v}$. MemVR enhances Mutual Information (MI) between $\hat{H}_{vq}$ and $Z_{v}$: $I(\hat{H}_{vq};Z_{v})\geq I(H_{vq};Z_{v})$.
The reinjection of $Z_{v}$ at intermediate layers of the model facilitates the replenishment of critical visual information, which may have been lost or distorted through earlier layers. This process increases MI between the hidden states and the visual features, ensuring adequate visual context is preserved.
Theorem 3.2.
Let $Y$ be the target output dependent on hidden states. If MI between ${H}_{vq}$ and ${Z}_{v}$ increases, then conditional entropy $H(Y\mid{H}^{(l)}_{vq})$ decreases with $H(Y\mid\hat{H}_{vq})\leq H(Y\mid H_{vq})$.
According to Theorem 3.1, 3.2, and DPI (Cover et al., 1991), MemVR improves the quality of the representations $H_{vq}$ and reduces the uncertainty in the output $Y$. As a result, the probability of hallucinations decreases, which is consistent with the observed findings of hallucination mitigation.
Theorem 3.3.
Within the Information Bottleneck (IB) framework, the loss of objective function, represented by the notation $\mathcal{L}(T)$, is optimized by MemVR, which is defined as $\mathcal{L}(\hat{H}_{vq})\leq\mathcal{L}(H_{vq})$, where $\mathcal{L}(H_{vq})=I(H_{vq};X_{vq})-\beta I(H_{vq};Y)$ is IB loss, and $\ beta$ is a trade-off parameter.
Anchored in the information bottleneck framework, MemVR optimizes the delicate balance between retaining relevant information from multimodal inputs and compressing non-essential details, thereby safeguarding the predictive performance of the hidden representations for the target output $Y$.
The theoretical underpinnings of MemVR are supported by the Data Processing Inequality (Cover et al., 1991) and the contraction properties of stochastic mappings in deep neural networks, as shown in various studies on the Information Bottleneck Principle (Achille & Soatto, 2018). By enhancing mutual information and reducing the uncertainty in hidden states, MemVR effectively mitigates hallucinations while preserving computational efficiency. Detailed proofs are in Appendix B.
Method   MSCOCO   A-OKVQA   GQA
%Accuracy %F1 Score %Accuracy %F1 Score %Accuracy %F1 Score
LLaVA1.5-7B 81.38  $\uparrow$0.0 79.65  $\uparrow$0.0 79.13  $\uparrow$0.0 79.10  $\uparrow$0.0 79.00  $\uparrow$0.0 79.13  $\uparrow$0.0
+ VCD (Leng et al., 2024) 84.66  $\uparrow$3.3 84.52  $\uparrow$4.9 80.99  $\uparrow$1.8 82.30  $\uparrow$3.2 81.74  $\uparrow$2.7 82.16  $\uparrow$3.0
+ OPERA (Huang et al., 2024a) 84.77  $\uparrow$3.4 85.46  $\uparrow$5.8 84.27  $\uparrow$5.1 84.08  $\uparrow$5.0 84.03  $\uparrow$5.0 83.83  $\uparrow$4.7
+ MemVR (Ours) 87.00  $\uparrow$5.7 85.87  $\uparrow$6.2 86.21  $\uparrow$7.0 86.64  $\uparrow$7.5 85.25  $\uparrow$6.2 85.59  $\uparrow$6.4
Qwen-VL-10B 83.79  $\uparrow$0.0 81.13  $\uparrow$0.0 84.74  $\uparrow$0.0 83.27  $\uparrow$0.0 84.41  $\uparrow$0.0 82.66  $\uparrow$0.0
+ VCD (Leng et al., 2024) 84.27  $\uparrow$0.4 82.12  $\uparrow$1.0 84.09  $\downarrow$0.7 82.53  $\downarrow$0.7 83.73  $\downarrow$0.7 82.75  $\uparrow$0.1
+ OPERA (Huang et al., 2024a) 84.93  $\uparrow$1.1 83.41  $\uparrow$2.3 - - - -
+ MemVR (Ours) 84.07  $\uparrow$0.3 81.55  $\uparrow$0.4 86.43  $\uparrow$1.8 85.56  $\uparrow$2.3 85.69  $\uparrow$1.3 84.53  $\uparrow$1.9
Table 2: Performance evaluation on POPE. The best results in each scenario are bolded for clarity. We report the averages under the three settings, e.g., Random, Popular, and Adversarial to show the robustness of the different methods directly. Green denotes improvement, and Red means degradation.
Method Commonsense QA(Reasoning)   Object-level Hallucination   Attribute-level Hallucination  Total Scores
Existence Count Position Color
LLaVA1.5-7B 110.71  $\uparrow$0.0 175.67  $\uparrow$0.0 124.67  $\uparrow$0.0 114.00  $\uparrow$0.0 151.00  $\uparrow$0.0 676.05
+ VCD (Leng et al., 2024) 112.86  $\uparrow$9.9 184.66  $\uparrow$9.0 138.33  $\uparrow$13.6 128.67  $\uparrow$14.6 153.00  $\uparrow$2.0 717.52
+ OPERA (Huang et al., 2024a) 115.71  $\uparrow$5.5 180.67  $\uparrow$5.0 133.33  $\uparrow$8.6 123.33  $\uparrow$9.3 155.00  $\uparrow$4.0 708.04
+ MemVR (Ours) 121.42  $\uparrow$18.5 190.00  $\uparrow$14.3 155.00  $\uparrow$30.3 133.33  $\uparrow$19.3 170.00  $\uparrow$19.0 769.75
Qwen-VL-10B 106.40  $\uparrow$0.0 155.00  $\uparrow$0.0 127.67  $\uparrow$0.0 131.67  $\uparrow$0.0 173.00  $\uparrow$0.0 693.74
+ VCD (Leng et al., 2024) 104.33  $\downarrow$2.1 156.00  $\uparrow$1.0 131.00  $\uparrow$3.3 128.00  $\downarrow$3.6 181.67  $\uparrow$8.6 701.00
+ OPERA (Huang et al., 2024a) 104.33  $\uparrow$2.2 165.00  $\uparrow$6.9 145.00  $\uparrow$4.8 133.33  $\uparrow$1.6 180.00  $\uparrow$7.0 727.66
+ MemVR (Ours) 120.00  $\uparrow$13.6 185.00  $\uparrow$30.0 145.00  $\uparrow$17.3 123.33  $\downarrow$8.3 185.00  $\uparrow$12.0 758.33
Table 3: Results on the hallucination subset of MME (including commonsense reasoning, existence, count, position, color scores). The best are in bold. MemVR achieves dramatic improvements.
4 Experiments
This section details the evaluation of our MemVR across three MLLMs on seven benchmarks.
4.1 Experiment Setup
Datasets and Metrics. To rigorously assess the effectiveness of our proposed method, we conduct a comprehensive set of experiments across two benchmarks specifically designed to evaluate hallucination mitigation and five general-purpose benchmarks to gauge the general performance:
  • Hallucination benchmarks: Polling-based Object Probing Evaluation (POPE) (Li et al., 2023c), and Caption Hallucination Assessment with Image Relevance (CHAIR) (Rohrbach et al., 2018);
  • General-purpose benchmarks: VizWiz-VQA (Gurari et al., 2018), MLLM Comprehensive Evaluation (MME) (Fu et al., 2023), Multimodal Benchmark (MMBench) (Liu et al., 2023b), Multimodal Veterinarian (MM-Vet) (Yu et al., 2024c), LLaVA-Bench (in-the-wild) (Liu et al., 2024).
More detailed information on these various benchmarks can be obtained from the Appendix C.1.
Backbones and Baselines. To evaluate our method, we utilize three well-known MLLMs: LLaVA-1.5 (Liu et al., 2024), Qwen-VL (Bai et al., 2023), and GLM4V (Wang et al., 2023). Further, We compare our methods with classic training-tree SOTA methods designed to mitigate object hallucination, including visual contrastive decoding SOTA VCD Leng et al. (2024), OPERA (Huang et al., 2024a) based on overconfidence penalty and hindsight allocation. As Dola (Chuang et al., 2023) is layer-wise contrastive decoding for LLMs and performs poorly in MLLMs, it will not be shown in the experiment. Experimental results are obtained and benchmarked using unified implementation.
Implementation Details. Greedy search is used as the default decoding strategy in MemVR for all benchmarks. For benchmarks, annotation questions are adapted to MLLM templates. For POPE, COCO, A-OKVQA, and GQA are used, while MMBench_DEV_EN is used for MMBench. MM-Vet is assessed using MM-Vet Online Evaluator, and gpt4-1106-preview is used for LLaVA-Bench. CHAIR uses images from COCO Val2014 with the query ”Please describe this image in detail”. In MemVR, do_sample=False, temperature=0, threshold=0.75, beam=1. All settings of the compared method follow the default configurations from the original papers. More details are in Appendix C.1.
Figure 5: Results on MMBench. MemVR enhances comprehensive performance on diverse tasks.
Strategy Static-3 Static-7 Static-11 Static-15 Static-19 Static-23 Static-27 Static-31 Dynamic
Preception 1508.08 1500.48 1526.40 1529.07 1513.96 1500.29 1510.47 1510.34 1512.80
Cognition 364.64 347.14 362.86 352.14 350.36 357.86 355.71 346.43 383.92
Total Score 1872.72 1847.62 1889.26 1881.22 1864.32 1858.15 1866.18 1856.77 1896.72
Table 4: Results of MemVR-static and MemVR-dynamic on MME. Static-# indicates fixation on layer # for visual retracing. MemVR-dynamic achieves optimal performance improvements.
4.2 Quantitative Results
Key Finding: MemVR consistently outperforms the baselines in mitigating hallucinations and improving overall accuracy across various scenarios.
Figure 6: Comparison of different injection ratios $\alpha$ on cognition scores.
Q1: How does MemVR performance on hallucination benchmark?
We conduct POPE, CHAIR, and MME evaluations, as shown in Table 2, Table 3 and Table 5, our MemVR obviously surpasses all of the compared baselines. For the results of POPE evaluation, we observe that our proposed MemVR presents robust effects. The performance of MemVR surpasses the baseline results by large margins, i.e., average up to +7.0% accuracy and +7.5% F1 score on A-OKVQA dataset under Random, Popular, and Adversarial settings. As showcased in Table 5, our MemVR achieves up to 15.6% improvements on CHAIRI compared with vanilla LLaVA-1.5-7B. For MME subset evaluations (encompass both object-level and attribute-level hallucinations), results in Table 3 show that MemVR achieves a uniform improvement in handling object-level and attribute-level hallucinations, as well as commonsense reasoning. Existence, Count, and Color scores all achieve dramatic improvements (Existence score in Qwen-VL up 30). On the contrary, Position scores are relatively low, which suggests weak position reasoning capability in MLLMs.
Model Length CHAIRS $\downarrow$ CHAIRI $\downarrow$ Recall $\uparrow$
LLaVA-1.5 100.6 50.0  $\uparrow$0.0 15.4  $\uparrow$0.0 77.1  $\uparrow$0.0
+ VCD 100.4 48.6  $\downarrow$1.4 14.9  $\downarrow$0.5 77.3  $\uparrow$0.2
+ OPERA 98.6 47.8  $\downarrow$2.2 14.6  $\downarrow$0.8 76.8  $\downarrow$0.3
+ MemVR 99.6 46.6  $\downarrow$3.4 13.0  $\downarrow$2.4 80.8  $\uparrow$3.7
Table 5: CHAIR hallucination evaluation results of LLaVA. Small values correspond to fewer hallucinations.
Q2: How well does MemVR perform on general-purpose benchmarks?
We evaluate the performance of MemVR on general-purpose benchmarks, i.e., VizWiz, MME, MMBench, MM-Vet, and LLaVA-Bench. Appx. C summarizes the results on MMBench, highlighting MemVR’s comparative performance relative to SOTA methods. As shown in Table 6, MemVR consistently outperforms competing models. Besides, MemVR achieves a significant improvement in overall performance listed in Table 7, with an average increase of 6.1% in OCR and spatial awareness tasks, demonstrating superior generalization capabilities. These results indicate that compared with CD-based methods, MemVR excels in hallucination mitigation and delivers competitive performance on general-purpose benchmarks. More complete results are in Appendix C.
Model Convs $\uparrow$ Detail $\uparrow$ Complex $\uparrow$ All $\uparrow$
LLaVA-1.5 58.8  $\uparrow$0.0 52.1  $\uparrow$0.0 74.6  $\uparrow$0.0 63.4  $\uparrow$0.0
+ VCD 57.8  $\downarrow$1.0 50.8  $\downarrow$1.3 77.9  $\uparrow$3.3 59.1  $\downarrow$4.3
+ OPERA 59.5  $\uparrow$0.7 49.6  $\downarrow$2.5 78.6  $\uparrow$4.0 59.8  $\downarrow$3.6
+ MemVR 63.8  $\uparrow$0.5 52.6  $\uparrow$0.5 77.9  $\uparrow$3.3 64.0  $\uparrow$0.6
Table 6: LLaVA-Bench evaluation results.
Q3: How efficient is MemVR?
MemVR operates dynamically based on the uncertainty, which employs visual retracing when the uncertainty exceeds threshold $\gamma$ on the early layer. If the uncertainty remains low across all layers—indicating that the model is highly confident in its generated results—MemVR is not triggered. This mechanism ensures efficient inference without extra computational overhead. Compared with VCD and OPERA, they need inference twice or the rollback strategy leads to exponentially added overheads, our MemVR only once regular inference.
Figure 7: MME evaluation results under different mixing ratios of noise and visual features.
Q4: How does visual uncertainty affect MemVR?
To examine the impact of visual uncertainty on MemVR ’s performance, we introduced Gaussian noise (McHutchon & Rasmussen, 2011) into the extracted visual features. We gradually increased the noise level to assess how MemVR ’s performance would respond, using MME as the benchmark and LLaVA as the backbone model. As illustrated in Figure 7, both the perception and cognition scores declined as the noise step increased. At high noise, the performance fell significantly below the baseline. This demonstrates that MemVR is sensitive to the quality of visual features and is efficient at understanding shallow information.
Figure 8: Results of GLM4V-9B. MemVR enhances comprehensive performance on diverse tasks.
Method gn_kw_rec rec ocr_sp ocr ocr_sp_rec ocr_kw_rec ocr_gn_sp Total
LLaVA1.5-7B 18.1  $\uparrow$0.0 67.6  $\uparrow$0.0 17.7  $\uparrow$0.0 48.3  $\uparrow$0.0 60.0  $\uparrow$0.0 21.2  $\uparrow$0.0 10.0  $\uparrow$0.0 31.1  $\uparrow$0.0
+ VCD 19.2  $\uparrow$1.1 62.2  $\downarrow$5.4 15.8  $\downarrow$1.9 29.2  $\downarrow$20.0 42.5  $\downarrow$17.5 17.5  $\downarrow$3.7 60.0  $\uparrow$50.0 30.2  $\downarrow$1.1
+ OPERA 21.8  $\uparrow$3.7 61.9  $\downarrow$5.7 21.5  $\uparrow$3.8 51.7  $\uparrow$3.4 56.2  $\downarrow$3.8 11.2  $\downarrow$10.0 30.0  $\uparrow$1.4 32.0  $\uparrow$0.9
+ MemVR 19.5  $\uparrow$1.4 70.3  $\uparrow$2.7 23.8  $\uparrow$6.1 48.3  $\uparrow$0.0 58.8  $\downarrow$1.2 21.2  $\uparrow$0.0 30.0  $\uparrow$20.0 32.4  $\uparrow$1.3
Table 7: MM-Vet evaluation results with multiple complicated multimodal tasks, where gn denotes language generation, kw means knowledge, sp denotes spatial awareness, and rec is recognition.
Figure 9: Visualisation of uncertainty across layers without and with MemVR. MemVR effectively reduces uncertainty after 8th layer, contributing to hallucination mitigations as Assumption 3.1.
Q5: How does layer uncertainty assist with visual retracing?
MemVR leverages layer entropy to trigger visual retracing dynamically. We compared the performance of fixed retracing layers with dynamic retracing layers and found that dynamic retracing outperformed fixed retracing, as demonstrated in Table 4. Under all experimental conditions, dynamic retracing achieved the highest total score on the MME evaluation using LLaVA. Furthermore, we analyzed the injection ratio for MemVR, with the optimal injection ratio observed around 0.15, as shown in Table 4. This indicates that a balanced level of $\alpha$ is crucial for maximizing performance.
Q6: Can MemVR adapt to other MLLMs?
Definitely, MemVR is designed to be flexible and compatible with various architectures. A key advantage of MemVR lies in its ability to function without requiring modifications to the Transformers library, facilitating smooth integration into both older and cutting-edge models. We conducted extensive experiments across multiple benchmarks with LLaVA, Qwen-VL, and GLM-4V, and all three models outperformed baselines.
4.3 Qualitative Study: What types of faults can our method address?
Figure 10: A case study in long text generation. MemVR effectively mitigates hallucinations.
In addition to evaluating single-word question-answering (QA) benchmarks, we further explore the models’ ability to generate comprehensive long-text descriptions for various tasks. As depicted in Figure 10, our MemVR excels in accurately identifying question-relevant details within images. In contrast, as discussed in Appendix C, Qwen-VL-Chat exhibits occasional difficulties in producing detailed image descriptions when applied to VCD. These shortcomings become evident in scenarios requiring a nuanced interpretation of image content. This suggests that MemVR demonstrates superior adaptability across different architectures, enabling more reliable long-text generation.
4.4 Limitations and Further Discussions
While MemVR demonstrates significant promise in improving the performance in MLLMs, it is not without limitations. One major challenge lies in the variability of MLLM architectures. Different MLLMs employ varying FNNs, activation functions, and knowledge systems, making it difficult to identify the optimal hyperparameters—such as injection ratios $\alpha$, and premature layer for each specific model. This requires considerable effort in model-specific tuning, which may limit the scalability of MemVR without further automation or standardization in hyperparameter selection.
Additionally, although our research focused on MLLMs with image inputs, MemVR is theoretically applicable to LLMs with different modalities. However, we have yet to explore its performance on inputs such as voice, point clouds, 3D meshes, or video frames. This opens up an exciting avenue for future work, where we plan to extend MemVR ’s framework to these diverse input formats and assess its efficacy across a broader range of tasks. Furthermore, we will explore integrating MemVR into the training procedures of MLLMs, rather than limiting its application to inference, to evaluate whether this could lead to even greater improvements in model performance and generalization.
5 Conclusion
This paper proposes a novel training-free paradigm to mitigate hallucination, named MemVR. In contrast to previous CD-based methods, which primarily focus on eliminating biases of language priors, MemVR seeks to replenish question-relevant visual clues towards more evidential answers, which signifies the other side of the coin. Our experiments, conducted on seven benchmarks, demonstrate the effectiveness of MemVR in mitigating hallucination and improving general performance.
    A: Ethic Considerations
  • Appendix
    B: Theoretical Analysis of MemVR in MLLMs
  • Appendix
    C: Additional Experiments, Results, and Discussions
Appendix A Ethic Considerations
We list some key ethical considerations of our method:
Bias and fairness. By injecting visual features from CLIP into the Feed-Forward Network (FFN) layers of LLMs, there’s a potential for inherited biases from the original models. CLIP, like many pre-trained models, may contain biases in how it represents certain objects, scenes, or demographics. These biases could propagate, affecting performance on different types of data (e.g., gender, ethnicity, cultural contexts). It’s important to evaluate how MemVR performs across diverse datasets and ensure that it doesn’t reinforce harmful stereotypes or disproportionately fail for certain groups.
Misuse of Enhanced Models. As MemVR aims to improve the long-text generation and overall performance of VLMs, enhanced models could be misused to generate deceptive or misleading content, such as deepfakes or disinformation. It’s important to consider whether there are safeguards in place to prevent malicious use of these improved models in scenarios like automated misinformation campaigns or unethical surveillance.
Data Privacy. If the benchmarks used for evaluating MemVR include datasets with personally identifiable information or sensitive content, care should be taken to ensure data privacy. Models should be evaluated on publicly available, anonymized, or ethically sourced datasets to avoid violating privacy laws or ethical norms.
Appendix B Theoretical Analysis of MemVR in MLLMs
In Multimodal Large Language Models (MLLMs), hallucinations often arise due to insufficient alignment between visual inputs and the model’s internal representations. This paper provides a rigorous theoretical analysis demonstrating that re-injecting visual features into the intermediate layers of MLLMs mitigates hallucinations and enhances representation capability.
We demonstrate that MemVR increases Mutual Information (MI) between the hidden states and visual tokens, decreasing the conditional entropy of outputs given the hidden state for fidelity to the visual input. We begin by defining the relevant variables and information-theoretic concepts that will be used throughout the proof as, $X_{vq}$ denote concatenated tokens of text and vision, with probability distribution $p({X}_{vq})$; ${Z}_{v}$ means visual (image) features, with probability distribution $p({Z}_{v})$; The output hidden states of the Transformer model at layer $k$, defined recursively as: $H_{vq}^{(k)}=f^{(k)}({H}_{vq}^{(k-1)},\mathbf{1}_{k=m}Z_{v})$, where $\mathbf{1}_{k=m}$ is the indicator function that equals 1 when $k=m$ (the layer where ${Z}_{v}$ is rejected) and 0 otherwise, and $Y$ denotes the target output of MLLMs.
The probability of hallucination can be expressed as:
$P_{\text{hallucination}}=P(Y\neq Y^{*}\mid X_{vq}),$
where $Y^{*}$ is the ground truth output. According to information theory, a higher conditional entropy $H(Y\mid X_{vq})$ indicates greater uncertainty of $Y$ given $X_{vq}$, which increases the probability of hallucination.
Information Flow of Visual Features. In a standard Transformer model, the initial input $X_{vq}$ undergoes multiple layers of processing. As the number of layers increases, the initial visual information may gradually diminish (information forgetting). In the absence of MemVR, the MI between the hidden states and the visual features $Z_{v}$ tends to decrease with depth:
$I(H_{vq}^{(l)};Z_{v})\leq I(H_{vq}^{(l-1)};Z_{v}),$
for $l>1$. This inequality indicates that in deeper layers, $H_{vq}^{(l)}$ contains less vision-related information.
Theorem B.1.
Assume that each Transformer layer acts as a deterministic or stochastic mapping with the Markov property. Then, the mutual information between the hidden states and the visual features decreases with depth:
$I(H_{vq}^{(l)};Z_{v})\leq I(H_{vq}^{(l-1)};Z_{v}).$
Proof.
Each Transformer layer can be modeled as a stochastic mapping (Markov kernel) that processes the input hidden states. Specifically, $H_{vq}^{(l)}$ is a function of $H_{vq}^{(l-1)}$, possibly incorporating additional inputs such as $Z_{v}$ at specific layers.
According to the Data Processing Inequality (DPI) (Cover et al., 1991), if $A\rightarrow B\rightarrow C$ forms a Markov chain, then:
$I(A;C)\leq I(A;B).$
In this context, consider $A=Z_{v}$, $B=H_{vq}^{(l-1)}$, and $C=H_{vq}^{(l)}$. Since $H_{vq}^{(l)}$ is generated from $H_{vq}^{(l-1)}$ without direct access to $Z_{v}$, we have the Markov chain $Z_{v}\rightarrow H_{vq}^{(l-1)}\rightarrow H_{vq}^{(l)}$. Applying DPI yields:
$I(Z_{v};H_{vq}^{(l)})\leq I(Z_{v};H_{vq}^{(l-1)}).$
Thus, mutual information between the hidden states and the visual features does not increase with depth. ∎
Visual Retracing in MLLMs. We reinject vision tokens $Z_{v}$ on $l$-th layer ($ahead\_layer\leq l<L$):
$\hat{H}_{vq}^{(l)}=\operatorname{FFN}_{\circlearrowright}^{(l)}(H_{vq}^{(l)}\propto Z_{v}).$
MemVR ensures that after the $l$-th layer, $\hat{H}_{vq}^{(l)}$ again contains question-aligned visual information.
Theorem B.2.
Let ${H}_{vq}$ be the hidden states of FFN and $\hat{H}_{vq}$ be after reinjection of visual evidence $Z_{v}$. MemVR enhances Mutual Information (MI) between $\hat{H}_{vq}$ and $Z_{v}$: $I(\hat{H}_{vq};Z_{v})\geq I(H_{vq};Z_{v})$.
Proof.
We aim to show that reinjecting $Z_{v}$ at layer $l$ increases the mutual information between the hidden states and $Z_{v}$ conditioned on $X_{vq}$.
By the definition of conditional mutual information:
$I(\hat{H}_{vq}^{(l)};Z_{v}\mid X_{vq})=\mathbb{E}_{X_{vq}}[I(\hat{H}_{vq}^{(l)};Z_{v}\mid X_{vq}=x)].$
Similarly,
$I(H_{vq}^{(l)};Z_{v}\mid X_{vq})=\mathbb{E}_{X_{vq}}[I(H_{vq}^{(l)};Z_{v}\mid X_{vq}=x)].$
Given $\hat{H}_{vq}^{(l)}=\operatorname{FFN}^{(l)}_{\Diamond}(H_{vq}^{(l)}\propto Z_{v})$ denotes the hidden states after utilizing MemVR on $l$-th, reinjection of $Z_{v}$ introduces a direct dependency between $\hat{H}_{vq}^{(l)}$ and $Z_{v}$ beyond what is present in $H_{vq}^{(l)}$. Since $\operatorname{FFN}^{(l)}_{\Diamond}$ is a deterministic function that incorporates $Z_{v}$, the mutual information $I(\hat{H}_{vq}^{(l)};Z_{v}\mid X_{vq})$ is at least as large as $I(H_{vq}^{(l)};Z_{v}\mid X_{vq})$. $\hat{H}_{vq}^{(l)}$ retains all information in $H_{vq}^{(l)}$ and additionally incorporates information from $Z_{v}$. Thus, MemVR ensures that:
$I(\hat{H}_{vq}^{(l)};Z_{v}\mid X_{vq})\geq I(H_{vq}^{(l)};Z_{v}\mid X_{vq}).$
By directly incorporating $Z_{v}$ into the computation of $\hat{H}_{vq}^{(m)}$, MemVR ensures that the hidden states retain more information about the visual features relative to the original hidden states $H_{vq}^{(m)}$, thereby increasing $I(\hat{H}_{vq}^{(m)};Z_{v}\mid X_{vq})$, enhancing the representation capability and utilizing visual information. ∎
Theorem B.3.
Let $Y$ be the target output dependent on hidden states. If MI between ${H}_{vq}^{(l)}$ and ${Z}_{v}$ increases, then conditional entropy $H(Y\mid{H}^{(l)}_{vq})$ decreases, leading to a lower probability of hallucinations:
$H(Y\mid\hat{H}_{vq}^{(l)})\leq H(Y\mid H_{vq}^{(l)}).$
Proof.
We aim to show that an increase in mutual information between $\hat{H}_{vq}^{(l)}$ and $Z_{v}$ conditioned on $X_{vq}$ leads to a decrease in the conditional entropy $H(Y\mid\hat{H}_{vq}^{(l)})$. According to the definition of conditional entropy, we have,
$H(Y\mid\hat{H}_{vq}^{(l)})=H(Y)-I(Y;\hat{H}_{vq}^{(l)}),$
$H(Y\mid H_{vq}^{(l)})=H(Y)-I(Y;H_{vq}^{(l)}).$
From Theorem B.2: $\hat{H}_{vq}^{(l)}$ contains more information about $Z_{v}$, i.e., $I(\hat{H}_{vq}^{(l)};Z_{v})\geq I(H_{vq}^{(l)};Z_{v})$. There is $I(Y;\hat{H}_{vq}^{(l)})\propto I(\hat{H}_{vq}^{(l)};Z_{v})$, thus we have $I(\hat{H}_{vq}^{(l)};Y)\geq I(H_{vq}^{(l)};Y)$. Then, we assume a dependency between $Z_{v}$ and $Y$, i.e., $I(Z_{v};Y)>0$, and subtract the inequalities, have:
$H(Y\mid\hat{H}_{vq}^{(l)})$ $=H(Y)-I(Y;\hat{H}_{vq}^{(l)})$
$\leq H(Y)-I(Y;{H}_{vq}^{(l)})$
$=H(Y\mid H_{vq}^{(l)}).$
Thus, MemVR reduces the conditional uncertainty of the target output given the intermediate embedding, thereby mitigating the probability of hallucinations and improving the model’s predictive capability. ∎
Theorem B.4.
Within the Information Bottleneck (IB) framework, reinjecting $Z_{v}$ at layer $m$ optimizes the objective function:
$\mathcal{L}(\hat{H}_{vq}^{(m)})\leq\mathcal{L}(H_{vq}^{(m)}),$
where the IB objective is defined as:
$\mathcal{L}(H)=I(H;X_{vq})-\beta I(H;Y),$
and $\beta$ is a trade-off parameter.
Proof.
The Information Bottleneck (IB) objective aims to find a representation $H$ that maximizes the mutual information with the target $Y$ while minimizing the mutual information with the input $X_{vq}$. The optimization objectives before & after MemVR are as follows:
$\mathcal{L}$ $=I(H_{vq}^{(l)};X_{vq})-\beta I(H_{vq}^{(l)};Y),$
$\mathcal{L}_{\Diamond}$ $=I(\hat{H}_{vq}^{(l)};X_{vq},Z_{v})-\beta I(\hat{H}_{vq}^{(l)};Y),$
where $I(\hat{H}_{vq}^{(l)};X_{vq},Z_{v})=I(\hat{H}_{vq}^{(l)};X_{vq})+I(\hat{H}_{vq}^{(l)};Z_{v}\mid X_{vq})$. The gap in the objective function is:
$\Delta\mathcal{L}$ $=\mathcal{L}_{\Diamond}^{(l)}-\mathcal{L}^{(l)}$
$=[I(\hat{H}_{vq}^{(l)};X_{vq})+I(\hat{H}_{vq}^{(l)};Z_{v}\mid X_{vq})-\beta I(\hat{H}_{vq}^{(l)};Y)]-[I(H_{vq}^{(l)};X_{vq})-\beta I(H_{vq}^{(l)};Y)]$
$=[I(\hat{H}_{vq}^{(l)};X_{vq})-I(H_{vq}^{(l)};X_{vq})]+I(\hat{H}_{vq}^{(l)};Z_{v}\mid X_{vq})-\beta[I(\hat{H}_{vq}^{(l)};Y)-I(H_{vq}^{(l)};Y)].$
To ensure that $\mathcal{L}_{\Diamond}^{(m)}\leq\mathcal{L}^{(m)}$, we require: $\Delta\mathcal{L}\leq 0.$ We define the changes in mutual information. Let $\Delta I_{X}=I(H_{vq}^{(l)};X_{vq})-I(H_{vq}^{(l-1)};X_{vq})$, $\Delta I_{Y}=I(H_{vq}^{(l)};Y)-I(H_{vq}^{(l-1)};Y)$. Note that $I(H_{vq}^{(l)};Z_{v}\mid X_{vq})\geq 0$. For $\Delta I_{X}$, the change in mutual information between $H_{vq}^{(l)}$ and $X_{vq}$ depends on how much additional information from $Z_{v}$ affects the dependence on $X_{vq}$. We denote the maximum possible increase as $\Delta I_{X}^{\max}$. For $\Delta I_{Y}$, From Theorem B.2, $\Delta I_{Y}\geq 0$, and suppose we can establish a minimum increase $\Delta I_{Y}^{\min}>0$. $I(H_{vq}^{(l)};Z_{v}\mid X_{vq})$ represents supplement information about $Z_{v}$ in $H_{vq}^{(l)}$ that is not already explained by $X_{vq}$, and we denote this maximum as $I_{\max}^{Z\mid X}$.
To satisfy this inequality, choose $\beta$ such that:
$\Delta\mathcal{L}\leq 0\Rightarrow\beta\Delta I_{Y}\geq\Delta I_{X}+I(H_{vq}^{(l)};Z_{v}\mid X_{vq}).$ (4)
Upper Bound on $\Delta I_{X}$ and $I(H_{vq}^{(l)};Z_{v}\mid X_{vq})$ as $\Delta I_{X}\leq\Delta I_{X}^{\max}$, $I(H_{vq}^{(l)};Z_{v}\mid X_{vq})\leq I_{\max}^{Z\mid X}$. Lower Bound on $\Delta I_{Y}$ as: $\Delta I_{Y}\geq\Delta I_{Y}^{\min}>0$. Then, we derive the condition with error bounds, for $\Delta\mathcal{L}\leq 0$, it suffices that:
$\beta\Delta I_{Y}^{\min}\geq\Delta I_{X}^{\max}+I_{\max}^{Z\mid X}\Rightarrow\beta\geq\frac{\Delta I_{X}^{\max}+I_{\max}^{Z\mid X}}{\Delta I_{Y}^{\min}}.$ (5)
This condition provides a lower bound for $\beta$ to ensure that reinjecting $Z_{v}$ at layer $m$ decreases the IB objective function. By adhering to this condition, MemVR optimizes the IB objective, balancing the trade-off between the compression of input information and the preservation of relevant information for prediction. ∎
By reducing the IB objective function, the model focuses more on information relevant to predicting $Y$ while compressing irrelevant information. The enhanced mutual information with $Y$ reduces the likelihood of generating hallucinated outputs not supported by the visual input.
Error Bounds Provide Guarantees. The upper and lower bounds on mutual information changes ensure that, under specific conditions (e.g., the selection of $\beta$), theoretical improvement holds.
Estimating the Bounds.
  • $\Delta I_{Y}^{\min}$ requires knowledge of how much additional information about $Y$ is gained by reinjecting $Z_{v}$. It can be estimated based on the mutual information $I(Z_{v};Y)$ and the effectiveness of $H_{vq}^{(m)}$ in capturing information relevant to $Y$.
  • $\Delta I_{X}^{\max}$ can be bounded based on the capacity of $H_{vq}^{(m)}$ to represent $X_{vq}$. Specifically, it relates to how much additional information $H_{vq}^{(m)}$ can encode about $X_{vq}$ beyond what was already captured in $H_{vq}^{(m-1)}$.
  • $H(Z_{v})$ is bounded by the entropy of the visual features, as mutual information cannot exceed the entropy of $Z_{v}$.
Through detailed mathematical derivations and the inclusion of upper and lower error bounds, we have established that:
  1. (a)
    Increased Mutual Information: Reinjecting visual features at an intermediate layer increases the mutual information between the model’s embeddings and the visual input.
  2. (b)
    Reduced Conditional Entropy: MemVR reduces the conditional uncertainty of the target output given the intermediate embedding, enhancing the model’s predictive accuracy and mitigating hallucination phenomena caused by the forgetting of visual information.
  3. (c)
    Optimization within IB Framework: Within the Information Bottleneck framework, MemVR optimizes the objective function, provided certain conditions on the mutual information changes are met and appropriate choices of the trade-off parameter $\beta$ are made.
These theoretical findings provide strong support for the practice of MemVR in MLLMs to improve their performance and reliability.
Appendix C Additional Experiments, Results, and Discussions
C.1 Benchmarks
In this appendix, we provide additional details into the benchmarks referenced in the main paper. To evaluate hallucinations, we employ the following five benchmarks:
CHAIR Rohrbach et al. (2018) evaluates how well the generated captions align with the content of the given image. CHAIR consists of two versions: CHAIR_S, which measures the inaccuracies at the sentence level, and CHAIR_I, which evaluates at the object level within the sentence by comparing the number of false objects to the total number of objects. For evaluation, we use the val2014 split of the MSCOCO Lin et al. (2014) dataset, which includes annotations for 80 object categories. We randomly select 500 images from the entire dataset and used the prompt “Please describe this image in detail.” for the MLLM.
Polling based Object Probing Evaluation (POPE) Li et al. (2023c) is a VQA-based metric proposed to assess hallucinations in MLLMs. This metric evaluates the MLLM’s response to the prompt “Is [object] is in this image?” To emphasize that this is a binary VQA task, we appended the prompt with “Please answer yes or no.” To select objects referenced in the question prompt, we followed three different sampling options: random, popular, and adversarial. We evaluated performance across all sampling options.
MLLM Evaluation (MME) Fu et al. (2023) evaluates the capabilities of MLLMs, dividing the evaluation into two major categories: perception and cognition. The perception category includes fine-grained tasks such as existence, count, location, rough color, poster, celebrity, scene, landmark, artwork identification, and OCR. The cognition category includes tasks like commonsense reasoning, numerical calculations, text translation, and code reasoning. All questions in this benchmark are structured to be answered with a simple yes or no.
Using the LLaVA-Bench Liu et al. (2024), we further demonstrated how well our proposed method maintains the language model performance. This benchmark involves posing various situational questions, such as dialogue, detailed descriptions, and complex reasoning, to randomly selected images from the MSCOCO val2014 dataset. A total of 60 questions are used to assess whether the model faithfully follows the instructions. The generated answers are evaluated by comparing them to the responses of a text-only GPT-4 model.
C.2 Reproducibility
Implementation details.
We employed greedy search as the default decoding strategy across all benchmark evaluations. For the hallucination benchmarks (POPE and CHAIR) and general-purpose benchmarks (MME, VizWiz-VQA, MMBench, MM-Vet, and LLaVA-Bench (in-the-wild)), questions from the annotation files were used as prompts, formatted to fit the chat templates of each respective MLLM. Specifically, we utilized the COCO, A-OKVQA, and GQA datasets for POPE evaluation, and MMBench_DEV_EN for MMBench. In the MM-Vet evaluation, we used an online evaluator powered by OpenAI GPT-4 to assess generated results, while for LLaVA-Bench (in-the-wild), we employed OpenAI’s model gpt4-1106-preview via API. For CHAIR, a randomly sampled image set from the COCO Val2014 dataset was used across all three models, with the prompt ”Please describe this image in detail.” We sampled three different sets of images using different random seeds and evaluated performance by calculating the mean and standard deviation of the results. All MemVR tests were conducted using a greedy decoding approach, with do_sample=False, temperature=0, threshold=0.75, and beam=1. For VCD tests, we set do_sample=True, temperature=1, noise_step=500, and the plausibility constraint hyperparameter $\lambda$ to 0.1, while $\alpha$, which controls the degree of contrastive emphasis, was set to 1, following the default parameter settings from the original code and literature. OPERA tests were configured with beam=5, sample=True, scale_factor=50, threshold=15, num_attn_candidates=5, and penalty_weights=1. Due to OPERA’s reliance on older versions of Torch and Transformers, it was incompatible with Qwen and GLM models, and thus experiments involving these models were not conducted. Additionally, our method introduces two hyperparameters: the informative layer l for activation calculations and the factor $\lambda$ to control the influence of entropy on the next token probability distribution. To map the hidden states from selected layers l to vocabulary tokens, we chose intermediate layers based on the model’s depth (e.g., layers 5 to 16 for vicuna-7b, which has 32 layers), and we set $\lambda$ as a fixed value (e.g., 0.75). All parameter settings adhered to the default configurations specified in the respective papers and code repositories.
Experimental Code.
To promote transparency and ensure the reproducibility of our work, we will release all experimental code, datasets, and detailed tutorials necessary for replicating our experiments. Our goal is to make it straightforward for researchers and practitioners to reproduce our results, regardless of their technical background. Additionally, by providing comprehensive documentation and clear guidelines, we aim to facilitate the extension of our method to other models and architectures, enabling the broader research community to explore its potential applications and improvements. We believe that open and reproducible research is essential for advancing the field and fostering collaboration.
Computational Resources.
Our experiments were conducted on eight A40 and four A800 GPUs. The computational bottleneck was not the numerical accuracy values but the collection of potential hallucinatory factors for analytical purposes, including logits and attention values for each head and layer.
C.3 Case Study
This case study aims to evaluate and present various benchmark cases across multiple domains systematically.
Figure 11: A case study comparing the levels of hallucination among various baselines
Figure 12: A case study comparing the levels of hallucination among various baselines
Figure 13: A case study comparing the levels of hallucination among various baselines
Figure 14: A case study comparing the levels of hallucination among various baselines
Figure 15: A case study comparing the levels of hallucination among various baselines
Figure 16: A case study comparing the levels of hallucination among various baselines
Figure 17: A case study comparing the levels of hallucination among various baselines
Figure 18: A case study comparing the levels of hallucination among various baselines
Figure 19: A case study comparing the levels of hallucination among various baselines
C.4 Additional Experiments and Results
Strategy 1-Token Len 5-Token Len 10-Token Len 20-Token Len 30-Token Len 50-Token Len 80-Token Len
Greedy 661.7 897.9 1273.1 1880.3 2501.8 3617.6 5256.6
Sample 786.8 1056.2 1314.9 1998.5 2568.5 3593.0 5587.0
VCD Sample 1747.74 2767.52 4027.07 4537.42 5031.39 7690.77 11569.3
Opera Beam 1566.1 3094.9 4166.4 6242.7 8436.9 12672.3 19247.2
MemVR Sample 750.8 1197.6 1780.5 2339.2 2631.7 3718.0 6011.0
MemVR Greedy 775.1 974.2 1337.5 1861.7 2742.8 4000.9 5545.5
Table A1: Time cost for generating tokens. All based on LLaVA1.5-7B
Method     LLaVABench (in-the-wild)
Average All_1 All_2 All_3
LLaVA1.5-7B 64.80  $\uparrow$0.0 63.40  $\uparrow$0.0 80.20  $\uparrow$0.0 50.80  $\uparrow$0.0
+ VCD (Leng et al., 2024) 63.20  $\downarrow$1.6 59.10  $\downarrow$4.3 82.00  $\uparrow$1.8 48.50  $\downarrow$2.3
+ OPERA (Huang et al., 2024a) 64.30  $\downarrow$0.5 59.80  $\downarrow$3.6 83.30  $\uparrow$3.1 49.80  $\downarrow$1.0
+ MemVR (Ours) 65.17  $\uparrow$0.4 64.00  $\uparrow$0.6 80.20  $\uparrow$0.0 51.30  $\uparrow$0.5
Qwen-VL-Chat 68.50  $\uparrow$0.0 70.40  $\uparrow$0.0 79.30  $\uparrow$0.0 55.80  $\uparrow$0.0
+ VCD (Leng et al., 2024) 53.77  $\downarrow$14.7 41.00  $\downarrow$29.4 85.30  $\uparrow$6.0 35.00  $\downarrow$20.8
+ OPERA (Huang et al., 2024a) - - - -
+ MemVR (Ours) 69.50  $\uparrow$1.0 69.50  $\downarrow$0.9 82.00  $\uparrow$2.7 57.00  $\uparrow$1.2
GLM-4V-9B 75.30  $\uparrow$0.0 88.40  $\uparrow$0.0 73.00  $\uparrow$0.0 64.50  $\uparrow$0.0
+ VCD (Leng et al., 2024) 74.23  $\downarrow$1.1 86.70  $\downarrow$1.7 72.80  $\downarrow$0.2 63.20  $\downarrow$1.3
+ OPERA (Huang et al., 2024a) - - - -
+ MemVR (Ours) 76.73  $\uparrow$1.4 88.90  $\uparrow$0.5 74.80  $\uparrow$1.8 66.50  $\uparrow$2.0
Table A2: Results on LLaVABench (in-the-wild) dataset. Best-performing method per model size and dataset is highlighted in bold; arrows indicate improvement or degradation over the baseline.
Method Total
LLaVA1.5-7B 31.1  $\uparrow$0.0
+ VCD (Leng et al., 2024) 30.20  $\downarrow$0.9
+ OPERA (Huang et al., 2024a) 32  $\uparrow$0.9
+ MemVR (Ours) 32.4  $\uparrow$1.3
Qwen-VL-Chat 49.0  $\uparrow$0.0
+ VCD (Leng et al., 2024) 34.60  $\downarrow$14.4
+ OPERA (Huang et al., 2024a) -
+ MemVR (Ours) 49.6  $\uparrow$0.6
GLM-4V-9B 63.4  $\uparrow$0.0
+ VCD (Leng et al., 2024) 59.40  $\downarrow$4.0
+ OPERA (Huang et al., 2024a) -
+ MemVR (Ours) 65.0  $\uparrow$1.6
Table A3: Results on MM-Vet dataset. Best-performing method per model size and dataset is highlighted in bold; arrows indicate improvement or degradation over the baseline, where lower values indicate better performance.
Method Accuracy
LLaVA1.5-7B 50.00  $\uparrow$0.0
+ VCD (Leng et al., 2024) 44.90  $\downarrow$5.1
+ OPERA (Huang et al., 2024a) 50.76  $\uparrow$0.8
+ MemVR (Ours) 51.50  $\uparrow$1.5
Qwen-VL-Chat 66.05  $\uparrow$0.0
+ VCD (Leng et al., 2024) 34.54  $\downarrow$31.5
+ OPERA (Huang et al., 2024a) -
+ MemVR (Ours) 66.36  $\uparrow$0.3
GLM-4V-9B 57.39  $\uparrow$0.0
+ VCD (Leng et al., 2024) 48.04  $\downarrow$9.4
+ OPERA (Huang et al., 2024a) -
+ MemVR (Ours) 58.00  $\uparrow$0.6
Table A4: Results on Vizwiz dataset. Best-performing method per model size and dataset is highlighted in bold; arrows indicate improvement or degradation over the baseline, where lower values indicate better performance.
Method     CHAIRS
Cs Ci Recall Len
LLaVA1.5-7B 47.60  $\uparrow$0.0 13.30  $\uparrow$0.0 80.60  $\uparrow$0.0 99.70  $\uparrow$0.0
+ VCD (Leng et al., 2024) 55.00  $\uparrow$7.4 15.80  $\uparrow$2.5 77.40  $\downarrow$3.2 101.20  $\uparrow$1.5
+ OPERA (Huang et al., 2024a) 47.60  $\uparrow$0.0 13.50  $\uparrow$0.2 79.00  $\downarrow$1.6 93.20  $\downarrow$6.5
+ MemVR (Ours) 46.60  $\downarrow$1.0 13.00  $\downarrow$0.3 80.80  $\uparrow$0.2 99.60  $\downarrow$0.1
GLM-4V-9B 40.40  $\uparrow$0.0 9.00  $\uparrow$0.0 72.70  $\uparrow$0.0 218.20  $\uparrow$0.0
+ VCD (Leng et al., 2024) 42.20  $\uparrow$1.8 9.60  $\uparrow$0.6 72.80  $\uparrow$0.1 239.80  $\uparrow$21.6
+ OPERA (Huang et al., 2024a) - - - -
+ MemVR (Ours) 39.40  $\downarrow$1.0 9.00  $\uparrow$0.0 70.70  $\downarrow$2.0 214.00  $\downarrow$4.2
Qwen-VL-10B 6.80  $\uparrow$0.0 5.30  $\uparrow$0.0 53.40  $\uparrow$0.0 17.60  $\uparrow$0.0
+ VCD (Leng et al., 2024) 13.00  $\uparrow$6.2 12.30  $\uparrow$7.0 47.90  $\downarrow$5.5 115.70  $\uparrow$98.1
+ OPERA (Huang et al., 2024a) - - - -
+ MemVR (Ours) 4.80  $\downarrow$2.0 3.30  $\downarrow$2.0 52.30  $\downarrow$1.1 15.00  $\downarrow$2.6
Table A5: Results on CHAIRS dataset. Best-performing method per model size and dataset is highlighted in bold; arrows indicate improvement or degradation over the baseline, where lower values indicate better performance.
Method     MMBench-Dev-EN
AR CP FP-C FP-S LR RR Overall
LLaVA1.5-7B 72.86  $\uparrow$0.0 75.68  $\uparrow$0.0 58.04  $\uparrow$0.0 63.48  $\uparrow$0.0 28.81  $\uparrow$0.0 51.30  $\uparrow$0.0 62.80  $\uparrow$0.0
+ VCD (Leng et al., 2024) 60.30 68.58 51.75 53.24 18.64 48.70 54.21
+ OPERA (Huang et al., 2024a) 69.85 75.00 56.64 66.21 28.81 53.04 62.80
+ MemVR (Ours) 71.86  $\uparrow$1.2 76.69  $\uparrow$1.0 57.34  $\downarrow$0.7 64.16  $\uparrow$0.9 31.36  $\uparrow$2.5 56.52  $\uparrow$5.2 63.75  $\uparrow$0.9
GLM-4V-9B 88.44  $\uparrow$0.0 86.49  $\uparrow$0.0 69.93  $\uparrow$0.0 85.67  $\uparrow$0.0 66.10  $\uparrow$0.0 85.22  $\uparrow$0.0 82.39  $\uparrow$0.0
+ VCD (Leng et al., 2024) 86.43 85.47 68.53 84.64 61.86 81.74 80.58
+ OPERA (Huang et al., 2024a) - - - - - - -
+ MemVR (Ours) 88.94  $\uparrow$0.5 86.49  $\uparrow$0.0 70.63  $\uparrow$0.7 86.01  $\uparrow$0.4 66.10  $\uparrow$0.0 85.22  $\uparrow$0.0 82.65  $\uparrow$0.3
Qwen-VL-10B 60.30  $\uparrow$0.0 71.28  $\uparrow$0.0 45.45  $\uparrow$0.0 62.80  $\uparrow$0.0 28.81  $\uparrow$0.0 38.26  $\uparrow$0.0 56.53  $\uparrow$0.0
+ VCD (Leng et al., 2024) 34.67 52.36 20.28 55.63 11.86 22.61 39.18
+ OPERA (Huang et al., 2024a) - - - - - - -
+ MemVR (Ours) 61.31  $\uparrow$1.0 71.28  $\uparrow$0.0 44.06  $\downarrow$1.4 62.80  $\uparrow$0.0 27.97  $\downarrow$0.8 38.26  $\uparrow$0.0 56.44  $\downarrow$0.1
Table A6: Results on MMBench dataset. Best-performing method per model size and dataset is highlighted in bold; arrows indicate improvement or degradation over the baseline, where lower values indicate better performance.
C.5 Supplement Implement Detail
The code of VCD Leng et al. (2024) is also released. However, the result of VCD evaluated in our experiments (e.g. POPE and MME benchmarks) is lower than the original paper. Therefore, we report the results in the original paper.
Appendix D Examples of capability integrations
Table A7: Six samples on MM-Vet benchmark requiring different capability integrations.
[Uncaptioned image] [Uncaptioned image]
Question: Which room is bigger, the double garage or the living room? Question: How many gallons of supreme gasoline can I get with \$50?
Ground Truth: Double garage Ground Truth: 13.6 $\mid$ 13.7
Required Capabilities: OCR, Spatial Awareness, Math Required capabilities: OCR, Math
[Uncaptioned image] [Uncaptioned image]
Question: Which car is on the parking spot 33? Question: Is this apple organic?
Ground Truth: No $\mid$ Empty Ground Truth: Yes
Required Capabilities: Recognition, OCR, Spatial Awareness Required capabilities: Recognition, OCR
[Uncaptioned image] [Uncaptioned image]
Question: What will the girl on the right write on the board? Question: Which are producers in this food web?
Ground Truth: 14 Ground Truth: Phytoplankton & Seaweed
Required capabilities: Recognition, OCR, Spatial Awareness, Math Required Capabilities: OCR, Knowledge, Spatial Awareness
D.1 Others
Q: How can the model understand information directly from the vision encoder, especially if it has a different vision system?
To ensure that MemVR is adaptable across diverse vision systems, we conducted experiments on multiple VLM architectures, including LLaVA, which utilizes a Visual-Instructional-Tuning framework with different sizes of ViT-based CLIP models, Qwen-VL-Chat, which employs a Q-Former-like architecture for visual processing, and ChatGLM-4v-9B, which integrates a large pre-trained visual encoder. These architectures encompass a broad range of vision models, providing confidence that MemVR is applicable to most VLMs in use today.
Artifacts and licenses
We report a list of licenses for all datasets and models used in our experiment in Table A8. We strictly follow all the model licenses and limit the scope of these models to academic research only.
Data Sources URL License
MSCOCO 2017 Link CC BY 4.0
ADE20K Link BSD-3-Clause
VQA Val Link CC BY 4.0
LLaVA-bench-in-the-wild Link Apache-2.0
ImageNet Link Custom License
MMBench Link Apache-2.0
Software Code URL License
LLaVA Link Llama Community Licence
Qwen-VL Link Tongyi Qianwen Licence
GLM-4V Link THUDM GLM-4 Licence
GPT-4V/4O Link OpenAI Term of Use
Table A8: License information for the scientific artifacts.

### v2

@@
 Abstract
-Despite their impressive capabilities, Multimodal Large Language Models (MLLMs) are susceptible to hallucinations, especially assertively fabricating content not present in the visual inputs.
-To address the aforementioned challenge, we follow a common cognitive process - when one’s initial memory of critical on-sight details fades, it is intuitive to look at them a second time to seek a factual and accurate answer.
-Therefore, we introduce Memory-space Visual Retracing (MemVR), a novel hallucination mitigation paradigm that without the need for external knowledge retrieval or additional fine-tuning.
-In particular, we treat visual prompts as supplementary evidence to be reinjected into MLLMs via Feed Forward Network (FFN) as “key-value memory”, when the model is uncertain or even amnesic about question-relevant visual memories.
-Comprehensive experimental evaluations demonstrate that MemVR significantly mitigates hallucination issues across various MLLMs and excels in general benchmarks without incurring added time overhead, thus emphasizing its potential for widespread applicability.
-[^1] Code can be available at https://github.com/1zhou-Wang/MemVR.
-614
-Figure 1: MemVR demonstrates strong performance across seven benchmarks spanning various domains (Left), with particularly outstanding results on the MME benchmark (Center).
-Additionally, MemVR is compared with contrastive decoding schemes, standing out for its ability to alleviate hallucinations using just a single inference, making it a more eco-friendly solution (Right).
+Despite their impressive capabilities, multimodal large language models (MLLMs) are prone to hallucinations, i.e., the generated content that is nonsensical or unfaithful to input sources.
+Unlike in LLMs, hallucinations in MLLMs often stem from the sensitivity of text decoder to visual tokens, leading to a phenomenon akin to “amnesia” about visual information.
+To address this issue, we propose MemVR, a novel decoding paradigm inspired by common cognition: when the memory of an image seen the moment before is forgotten, people will look at it again for factual answers.
+Following this principle, we treat visual tokens as supplementary evidence, re-injecting them into the MLLM through Feed Forward Network (FFN) as “key-value memory” at the middle trigger layer.
+This look-twice mechanism occurs when the model exhibits high uncertainty during inference, effectively enhancing factual alignment.
+Comprehensive experimental evaluations demonstrate that MemVR significantly mitigates hallucination across various MLLMs and excels in general benchmarks without incurring additional time overhead.
+The implementation is available from https://github.com/1zhou-Wang/MemVR.
 1 Introduction
-Multimodal Large Language Models (MLLMs), due to their formidable capacity to comprehend visual inputs, have emerged as indispensable tools in computer vision (Koh et al., 2024) and natural language processing (Tu et al., 2023) to tackle numerous visual tasks and facilitate complex visual question answering (VQA).
-Nevertheless, MLLMs still exhibit certain limitations, i.e., the so-called “hallucination” phenomenon (Huang et al., 2024b; Zheng et al., 2024).
-Specifically, MLLMs frequently generate descriptions inconsistent with user-provided visual inputs, such as incorrectly outputting non-existent things in the image or conflicting judgments.
-This flaw poses significant risks to the reliability of MLLMs as trustworthy assistants (Yu et al., 2024b), particularly in safety-critical applications (Zou et al., 2023) (e.g., clinical healthcare (Lin et al., 2024) and autonomous driving scenarios (Ding et al., 2024)).
-While the exact reasons for hallucinations in MLLMs are not fully understood, one possible factor could be the imbalance between their understanding of visual and textual information.
-Typically, MLLMs encode images into vision tokens via CLIP (Radford et al., 2021), which are fed along with text tokens into LLMs for decoding.
-LLMs excel at text comprehension but struggle with visual information perception and memory, where differences in information density between modalities may cause inconsistencies during decoding.
-Numerous methods have been proposed to mitigate hallucinations in MLLMs, and general studies can be broadly categorized into three streams: (i) Retrieval-Augmented Generation (RAG) (Shuster et al., 2021; Caffagni et al., 2024) that incorporates knowledge from external databases to mitigate the problem of “hallucination”, as well as (ii) through extra Fine-tuning (Yu et al., 2024b) to enhance the self-consistency of generation, and (iii) Contrastive Decoding (CD) strategy, which do not involve extra training.
+Multimodal Large Language Models (MLLMs), known for their ability to process visual, auditory and textual data, are crucial in fields such as computer vision (Koh et al., 2024) and natural language processing (Tu et al., 2023), helping with visual tasks and complex visual question answering.
+However, MLLMs still face challenges, notably the “hallucination” issue (Huang et al., 2024b; Zheng et al., 2024; Lyu et al., 2025), where they generate contents inconsistent with original inputs, such as generating nonexistent objects or conflicting judgments.
+This flaw undermines their reliability, especially in areas critical to safety such as healthcare (Lin et al., 2024) and autonomous driving (Ding et al., 2024).
+Although the causes of hallucinations are unclear, one potential factor is the imbalance between their understanding of visual and textual modalities.
+This imbalance may induce biases when the model integrates multimodal information, leading to outputs that do not match objective facts.
+Currently, numerous methods are being tried out to solve this problem.
+General studies can be broadly categorized into four streams: (i) Retrieval-Augmented Generation (RAG) (Qu et al., 2024) which incorporates knowledge from external databases to mitigate hallucinations, as well as (ii) through extra fine-tuning (Yu et al., 2024a) to enhance the self-consistency of generation; (iii) attention intervention (Huang et al., 2024a) and (iv) Contrastive Decoding (CD) (Leng et al., 2024) strategies, which not involve extra training.
 Specifically, RAG and fine-tuning patterns typically employ external knowledge retrieval or robust instruction-tuning datasets to post-hoc debias (Yang et al., 2024; Liu et al., 2023a), which inevitably introduces substantial computational overhead or storage requirements.
-For example, some approaches(Yin et al., 2023; Yu et al., 2024a) have fine-tuned models using high-quality visual instructions generated by advanced automated annotation tools including GPT-4 (OpenAI, 2023).
-CD-based methods (Li et al., 2023b; Shi et al., 2024) represent a simpler and more efficient way to mitigate hallucinations than RAG and fine-tuning based methods.
-Particularly, CD-based hallucination mitigation methods usually modulate the logits of the next token prediction through contrast manner or penalty mechanisms.
-As illustrated in Figure 2 (a), DoLa (Chuang et al., 2023) enriches factual knowledge via layer-wise contrasting and reduces the generation of incorrect facts in LLMs.
-In MLLMs, VCD (Leng et al., 2024) amplifies the language priors by adding Gaussian noise to the visual inputs, thereby reducing over-reliance on statistical biases and single-modal priors through contrasting output distributions from original and distorted visual inputs as in Figure 2 (b).
-This perturbation of original inputs requires task-specific design, inevitably doubling inference costs.
-More critically, contrastive distributions are agnostic to visual and instructional nuances, which may not always amplify the intended hallucinations, occasionally introducing potential noise into CD.
-Figure 2: The conventional paradigm for hallucination mitigation, employing contrastive decoding methods such as DoLa (a) and VCD (b), is compared to our proposed MemVR strategy (c).
-Method 20-Token Len 50-Token Len 80-Token Len
-LLaVA-1.5 1880.3  $\downarrow$×$\times$1.0 3617.6  $\downarrow$×$\times$1.0 5256.6  $\downarrow$×$\times$1.0
-+ VCD 4537.4  $\uparrow$×$\times$2.4 7690.8 $\uparrow$×$\times$2.1 11569.3 $\uparrow$×$\times$2.2
-+ OPERA 6242.7  $\uparrow$×$\times$3.3 12672.3  $\uparrow$×$\times$3.5 19247.2  $\uparrow$×$\times$3.7
-+ MemVR 1861.7  $\uparrow$×$\times$1.0 4000.9  $\uparrow$×$\times$1.1 5545.5  $\uparrow$×$\times$1.1
-Table 1: Efficiency Comparisons for generating different length tokens, using an NVIDIA-A40 GPU. Inference time (ms) of different methods is recorded.
-In this work, we delve into the challenges of hallucination mitigation in MLLMs and address the shortcomings of CD-based approaches.
-Our research is grounded in a common cognitive process: when the initial memory of certain critical visual details fades, it is intuitive to look at them for the second time to search for the accurate answer (O’regan, 1992; Ballard et al., 1995; Horowitz & Wolfe, 1998).
-Different from visual contrastive decoding strategies that alleviate hallucinations by diminishing language priors (Leng et al., 2024; Qu et al., 2024; Park et al., 2024), we propose a novel Memory-space Visual Retracing (MemVR) method that mitigates hallucinations through supplementing visual evidence, akin to the two sides of a coin.
-MemVR, as shown in Figure 2 (c), is an architecture-agnostic, plug-and-play solution that re-injects visual features into an intermediate layer suffering from vision-related memory lapse with only one regular inference.
-This novel hallucination mitigation paradigm in terms of efficiency, and its inference cost and performance are optimal compared to previous studies as listed Table 1.
-It’s a game-changer for effectiveness and efficiency.
-Through extensive experiments on multimodal hallucination benchmarks, as well as GPT-4o [^2] GPT-4o-2024-08-06: https://platform.openai.com/docs/models/gpt-4o.
-evaluations, we show the comprehensive performance improvements of MemVR in hallucination mitigation and general capabilities.
-Our contributions can be summarized as follows:
-  • We propose MemVR, a novel training-free hallucination mitigation paradigm that effectively alleviates hallucinations in MLLMs.
-In contrast to previous methods, which primarily focus on eliminating biases of language priors, MemVR seeks to replenish question-relevant visual clues towards more evidential responses, which signifies the other side of the coin.
-  • We design a dynamic premature layer injection strategy with visual retracing in MLLMs, mimicking human intuitive thinking to revisit image features for self-consistency and credible answers when pivotal memories are scrambled.
-Furthermore, we theoretically demonstrate that visual retracing can effectively diminish hallucinations from an information-theoretic perspective.
-  • Comprehensive experimental results demonstrate the effectiveness of MemVR in mitigating hallucinations and enhancing general cognitive and perceptual performance, as well as its high efficiency.
-Our research will make a substantial contribution to trustworthy multimodal intelligence.
+Attention intervention, though not requiring additional data, usually involves retrospection-allocation operations, which bring about high inference latency and a large memory footprint.
+Figure 1: Comparison of the conventional CD-based hallucination mitigation paradigm VCD, and our proposed efficient MemVR.
+CD-based methods (Li et al., 2023a; Shi et al., 2024) represent a simpler and more efficient way to mitigate hallucinations than other paradigms.
+Specifically, CD-based hallucination mitigation paradigms, represented by VCD (Leng et al., 2024), modulate logits of the next token prediction in a contrastive manner.
+As shown in Figure 1 left, VCD amplifies the language priors by adding Gaussian noise to the visual inputs, reducing over-reliance on statistical biases and single-modal priors through contrasting output distributions from original and distorted visual inputs.
+This perturbation of original inputs requires task-specific design, inevitably ✄ doubling inference costs.
+More critically, contrastive distributions are agnostic to visual and instructional nuances, which may not always amplify the intended hallucinations, occasionally ✄ introducing potential noise into CD.
+In this work, we delve into the challenges of hallucination mitigation in MLLMs and propose a novel paradigm that overcomes the two limitations of CD-based approaches.
+Our research is grounded in a common cognitive process: when the memory of an image seen the moment before is forgotten, it is intuitive to look twice for accurate and factual answers (Ballard et al., 1995; Horowitz & Wolfe, 1998).
+Following this principle, we propose Memory-space Visual Retracing (MemVR) that mitigates hallucinations through supplementing visual evidence, which can also be called look-twice mechanism.
+As shown in Figure 1 right, MemVR reinjects visual tokens through Visual Retracing (VR), i.e., look-twice mechanism, into the middle trigger layer suffering from high uncertainty, without incurring additional inference cost.
+Compared with VCD and other approaches, our proposed look-twice mechanism is optimal in terms of performance, efficiency, and memory cost as shown in Figure 2 and Table 1.
+Through extensive experiments on multimodal hallucination benchmarks, as well as GPT-4o evaluations, including eight public benchmarks, we show the comprehensive performance improvements of MemVR in hallucination mitigation and general capabilities.
+The main contributions can be summarized as follows:
+  • We propose MemVR, a novel, efficient, minimalist, and plug-and-play approach that achieves both model fidelity and efficiency, which reinforces attention to visual information for enhancing modality balance during the forward pass, without eliminating beneficial language priors.
+  • We present static and dynamic VR strategies that shift hidden states of the intermediate layer in MLLM for self-enhancement, rather than modulating logits directly in a CD manner, thus avoiding multi-round decoding.
+  • Our analysis reveals that hallucinations are triggered by the sensitivity of text decoder (i.e., LLM) to non-text modality. This finding is experimentally validated.
+  • Comprehensive experiments and evaluations on multiple models demonstrate that MemVR outperforms SOTA methods in performance and inference speed, such as +7.0% on the POPE benchmark, 15.6% improvement on CHAIRI, and total score of MME increased by +32.2 marks.
+The results on popular hallucination and general benchmarks validate the generalizability of our method.
+Figure 2: Radar charts comparing models across benchmarks.
+Method Latency $\downarrow$  (ms/token)    Throughput $\uparrow$   (token/ms) Time Cost $\downarrow$ for 80 Tokens (ms) Memory $\downarrow$ Cost (MB)
+Greedy 65.71 ($\times$1.00) 0.015 ($\times$1.00) 5256.6 ($\times$1.00) 14257 ($\times$1.00)
+Sample 69.84 ($\times$1.06) 0.015 ($\times$1.00) 5587.0 ($\times$1.00) 14262 ($\times$1.00)
+OPERA 240.59 ($\times$3.66) 0.004 ($\times$0.27) 19247.2 ($\times$3.66) 21300 ($\times$1.49)
+ICD 70.84 ($\times$1.08) 0.014 ($\times$0.93) 5666.9 ($\times$1.08) 14263 ($\times$1.00)
+VCD 144.62 ($\times$2.20) 0.007 ($\times$0.47) 11569.3 ($\times$2.20) 14967 ($\times$1.05)
+MemVR 68.32 ($\times$1.04) 0.015 ($\times$1.00) 5545.5 ($\times$1.06) 14345 ($\times$1.01)
+Table 1: Performance comparison of SOTA methods and MemVR in latency, throughput, time cost, and memory usage. The best and suboptimal results are highlighted in green and blue, respectively.
 2 Related Work
+Table 2: Comprehensive comparisons between the proposed MemVR and existing approaches are presented.
+MemVR introduces a low-latency look-twice decoding mechanism, optimizing hidden states to support multimodal integration and enhance overall performance.
+MemVR uniquely achieves visual hallucination mitigation and general improvement.
+SOTA methods we compare are emphasized in gray.
+Methods
+ VH
+   Mitigation
+General
+Improvement
+Inference
+Latency
+Expand to
+More Modalities
+Modified
+Component(s)
+Decoding Paradigm
+DoLa (Chuang et al., 2023) Negative Low Logits Contrastive
+OPERA (Huang et al., 2024a) Medium High Attention matrix Att-intervention
+EAH (Zhang et al., 2024) Medium High Attention matrix Att-intervention
+CCA (Xing et al., 2024) Medium High Attention matrix Att-intervention
+ICD (Wang et al., 2024) Medium Medium Textual input, logits Contrastive
+ID (Kim et al., 2024) Medium Medium Textual input, logits Contrastive
+SID (Huo et al., 2024) Medium Medium Visual tokens, logits Contrastive
+VCD (Leng et al., 2024) Low Medium Visual input, logits Contrastive
+HALC (Chen et al., 2024b) Medium Ultrahigh Visual input, logits Contrastive
+VORD (Neo & Chen, 2024) Medium Medium Visual input, logits Contrastive
+MemVR (ours) High Low Hidden states Look-twice
 MLLMs and Challenges.
@@
 Leveraging open-source Large Language Models (LLMs) such as LLaMA families (Touvron et al., 2023), MLLMs (Alayrac et al., 2022; Wu et al., 2024) have demonstrated enhanced adaptability across a range of visual language tasks, leading to a more profound ability to interpret the world.
-Models like LLaVA (Liu et al., 2024), Qwen-VL (Bai et al., 2023), and GLM4V (Wang et al., 2023) have further advanced this field, enabling users to interact with these agents using both image and text prompts.
+Models like LLaVA (Liu et al., 2024b), Qwen-VL (Bai et al., 2023), GLM4V (Wang et al., 2023), and LLaVA-Next (Liu et al., 2024a) have further advanced this field, enabling users to interact with these agents using both image and text prompts.
 These models adhere to two critical training phases: pre-training feature alignment and instruction fine-tuning, ensuring they better comprehend the format of instruction inputs (Yin et al., 2024).
-However, despite their impressive performance in many areas, MLLMs still suffer from hallucination issues.
-Thus, in this paper, we primarily conducted experiments and analysis on these three representative models.
-Hallucinations in MLLMs.
-Before LLMs, hallucination in NLP was mainly seen as generating nonsensical or deviant content.
-In MLLMs, “hallucination” is defined as the model generating content inconsistent with the provided image.
-This issue stems significantly from inadequate alignment among modalities.
-Several methods have been explored to mitigate hallucinations in MLLMs.
-Early efforts focused on fine-grained modality alignment (Rohrbach et al., 2018) and reducing co-occurrence biases (Kim et al., 2023) in small-scale VLMs, but these approaches struggle to scale with MLLMs.
-More recent strategies involve hallucination-targeted datasets for fine-tuning (Gunjal et al., 2024), post-hoc revisors (Zhou et al., 2024), and adopting RLHF (Yu et al., 2024b).
+However, despite their impressive performance in many areas, multimodal large language models still suffer from hallucination issues.
+In this work, we conducted experiments and analysis on these representative models to validate MemVR’s applicability.
+Mitigating Hallucinations in MLLMs.
+Researchers have made extensive efforts to uncover the causes of hallucinations (Yin et al., 2023; Zhou et al., 2023; Bai et al., 2024).
+Early works to mitigate hallucinations focused on fine-grained modality alignment (Rohrbach et al., 2018) and reducing co-occurrence biases (Kim et al., 2023) in small-scale models.
+More recent strategies involve hallucination-related datasets for fine-tuning (Gunjal et al., 2024), post-hoc revisors (Zhou et al., 2024), and adopting RLHF (Yu et al., 2024a).
+LURE (Zhou et al., 2023) trains a reviser to edit the possible hallucinated words in the responses.
 While effective, these methods are resource-intensive.
-CD-based approaches (Chuang et al., 2023; Leng et al., 2024) adjust the decoding distribution to mitigate hallucinations, but it does not consistently improve performance.
-Compared with them, our MemVR stands as “a paradigm of effectiveness and efficiency” in hallucination mitigation, effortlessly enhancing performance without extra training.
+Attentional intervention strategies (Zhang et al., 2024; Xing et al., 2024), represented by OPERA (Huang et al., 2024a), are simpler and do not require additional data or training, but have a higher inference latency.
+CD-based approaches (Chuang et al., 2023; Chen et al., 2024b; Neo & Chen, 2024), represented by ICD (Wang et al., 2024) and VCD (Leng et al., 2024), adjust the decoding distribution to mitigate hallucinations in MLLM, but this does not consistently improve performance as it introduces potential noise into the CD.
+Comparisons.
+Distinct from these methods, our proposed MemVR offers a novel decoding strategy that effectively reduces visual hallucinations (VH) without necessitating extra models, data, and training.
+Table 2 illustrates the differences and advantages of our method compared to recent representative SOTA approaches.
+DoLa (Chuang et al., 2023) targets hallucinations in LLM and is negative for alleviating VH, although it exhibits low latency.
+Attentional intervention methods (Zhang et al., 2024; Xing et al., 2024), represented by OPERA (Huang et al., 2024a), have achieved considerable progress in mitigating VH, but are limited to the problem of high latency.
+CD-based approaches, whether modifying the textual input (Wang et al., 2024; Kim et al., 2024) or visual input (Leng et al., 2024; Chen et al., 2024b; Neo & Chen, 2024), both aim to reduce the output probability of incorrect tokens through the comparison of logits.
+However, this brings two challenges, one is that the CD strategy may introduce noise to the output distribution, thus losing the original capability, and the other is that the CD-based mechanisms often require multiple rounds of inference to obtain several pairs of logits for contrasting, resulting in a high latency.
+Significantly, these methods work negatively or fail in general-purpose testing.
+Compared with them, MemVR stands as “a paradigm of effectiveness and efficiency” in visual hallucination mitigation and general improvement.
+3 Background and Motivation
+3.1 Problem Formulation
+Given an MLLM $\mathcal{M}_{\theta}^{\text{MLLM}}$ parameterized by $\theta$, with a general architecture consisting of a text embedding layer, a vision encoder, a vision-text interface module, a text decoder consisting of $L$ number of transformer layers, and an affine layer $\varsigma(\cdot)$ which predicts the distribution of the next token.
+For an image-grounded text generation task, given a textual query $x$ and an input image $v$, MLLM first extracts vision features of $v$ by the vision encoder, and then converts them into visual tokens $z_{v}$ by MLP or Q-Former (Wadekar et al., 2024) modules.
+Aligned vision tokens $z_{v}$ are concatenated with the query $x$ as input to the text decoder, and finally decoded into a textual response $y$ autoregressive, which is formulated as follows:
+$y_{t}\sim p_{\theta}(\cdot|v,x,y_{<t})\propto\text{softmax}(f_{\theta}(\cdot|v,x,y_{<t})),$ (1)
+where $y_{t}$ indicates the $t^{th}$ token, $y_{<t}$ is the token sequence generated up to time step $t$, and $f_{\theta}$ is the logit distribution, i.e., unnormalized log-probabilities produced by $\mathcal{M}_{\theta}^{\text{MLLM}}$.
+When the text generation $y$ is inconsistent or in conflict with the input image $v$, MLLM is believed to present hallucination issues.
+The objective of visual hallucination mitigation is to minimize the appearance of incorrect or conflicting tokens, ensure faithfulness to $v$ when answering the query $x$, and simultaneously maintain high-quality text generations.
+3.2 Why Does MLLM Generate Hallucination?
+Hallucinations in MLLM are caused by multiple factors, including inherent biases in the training data (Zhou et al., 2023), visual uncertainty resulting from the model’s statistical bias and priors (Leng et al., 2023), and the limitations of current models in accurately discerning context and fact throughout the output generation process (Daunhawer et al., 2021).
+Upon a more in-depth analysis, we consider that the imbalance of modalities in MLLM and the autoregressive characteristic of language models are likely crucial factors causing their hallucinatory phenomena.
+Taking image and text as an example, since an image possesses a much higher information density than a piece of text, it is reasonable to suppose that LLMs struggle to understand or memorize vision information compared to text.
+Moreover, autoregressive decoding causes MLLMs to increasingly depend on textual information, including query $x$ and growing tokens $y_{<t}$, inevitably decreasing reliance on visual input.
+As PAI(Liu et al., 2024c) proposed to mitigate hallucinations by paying more attention to images, this reflects the fact that hallucinations may be caused by modality imbalances.
+Figure 3: Uncertainty of different layers to predict the next token. Rows denote indices of the early layers, and column names are decoded tokens in each step. Uncertainty distribution is dynamic.
+Figure 4: (Left) Performance under different scaling ratios to text / image feature value on MME; (Right) Performance changes when look-twice to text / image / text+image (i.e, @t+i), respectively.
+To further verify this conjecture, we scale the feature values of different modalities up or down to simulate the modality imbalance phenomenon.
+As shown in the left of Figure 4, it can be observed that: when the image and text features are scaled up proportionally, modality balance is disrupted in both cases, yet the performance decline caused by magnifying the image features is more significant than that of text features; however, when the image and text features is scaled down proportionally, the contraction on text features led to a drastic performance drop, while the scaling down of visual features had a relatively minor impact.
+We can obtain three findings in this experiment: ① modality imbalance issues cause hallucinations in MLLMs, ② text decoders (i.e., LLMs) are more text-informed, ③ LLMs are harder to comprehend the visual modality than textual inputs.
+This modality imbalance leads to a substantial deviation from the accurate representation of visual input, eventually giving rise to hallucinations, as manifested by the phenomenons in the aforementioned studies (Zhou et al., 2023).
+This is especially evident in generating longer responses, explaining the correlation between higher VH and larger maximum token lengths, as presented in Huang et al.
+(2023).
+Figure 5: Uncertainty distribution across layers during token reasoning in hallucinations. Red-outlined regions show higher uncertainty in middle and late layers for hallucinatory tokens: a, pom.
+3.3 The Pattern of Hallucinations in MLLM
+In order to further explore the potential pattern of hallucinations in MLLM, this study employs uncertainty as the metric.
+As findings of Chen et al.
+(2024a) in LLMs: incorrect tokens generally exhibit higher entropy than correct ones, we also observe this phenomenon in MLLMs, the visualization case is shown in Figure 5.
+From Figure 5, it can be noticed that when the model generates illusion tokens, i.e., a, pom -eg-ran-ate (pomegranate), the uncertainty in the middle and last layers of the model is high in MLLM.
+Uncertainty quantification.
+Following the DoLa (Chuang et al., 2023), we compute the probability of the next token via the vocabulary head $\varsigma(\cdot)$ on each layer during reasoning.
+Then, we introduce an entropy-based metric (Farquhar et al., 2024) to quantify the output uncertainty of each layer in the text decoder as $u=\sum-p_{i}\log p_{i}/\log{N}$, where $\{p_{i}\}_{i=1}^{N}$.
+In addition, in the context of tokens involving objects, attributes or relations, uncertainty is also high.
+We conduct a preliminary analysis with 32-layer LLaVA-1.5-7B.
+Specifically, we compute the uncertainty in the output distributions of early layers.
+Figure 3 shows the uncertainty scores of different early layers when decoding the answer, we can observe that the computed uncertainty remains relatively high in later layers when predicting key entity objects, attributes, or relations, such as wooden, filled, foods.
+This phenomenon suggests that LLM is still uncertain about its predictions in the last few layers and may inject more factual knowledge into the predictions.
+On the other hand, when predicting function words and those tokens copied from the question, e.g., image, with, we observe that the uncertainty becomes very low from the middle layers.
+This finding implies that the model is deterministic for easy-to-predict tokens at the intermediate layer and keeps the distribution of outputs almost constant at higher layers, however, it is more uncertain for difficult-to-predict key tokens and may constantly change its predictions until the final layer.
+3.4 Refresh Visual Memory Alleviates Hallucination
+Based on the findings, we propose that both overreliance on textual information and schema imbalance contribute to a phenomenon akin to “amnesia” about visual information.
+It is easy to understand that from the shallow to the deep layers of the Transformer, attention is progressively biased towards textual tokens, resulting in visual tokens at the deeper layer that hardly affect the results, which is consistent with how the attention intervention strategies work.
+To address this, we try to refresh visual information for trustworthy answers when the model encounters high uncertainty.
+We verify this hypothesis through an empirical study.
+As shown in the right of Figure 4, the proposed look-twice strategy is used for combinations of different modalities, including text, image, text+image, where it can be found that the best performance improvement is achieved while only replenishing image information to the model, i.e., refreshing visual memory can effectively alleviate hallucinations in MLLMs.
+This study lends further credence to our conjecture regarding the causes of hallucinations in Sec.
+§3.2.
 4 Methodology
-In this section, we first formulate the generation process of MLLMs to facilitate a clearer understanding of our MemVR.
-Moreover, we introduce our hypothesis about the causes of hallucinations and discuss visual retracing and its dynamic strategy.
-Further, we conduct theoretical analysis.
-3.1 MLLMs Generation Process
-Typically, MLLMs are composed of a visual encoder ${C}_{v}$, a text embedding layer, $L$ number of transformer layers, and an affine layer $\zeta(\cdot)$ which predicts the distribution of the next token, with diverse modalities as inputs, e.g., images and text.
-Regardless of specific architectural variations, MLLMs commonly employ ${C}_{v}$ to extract vision tokens from the raw image and project it into modality-shared feature space via an MLP or Q-Former module (Wadekar et al., 2024).
-Aligned vision tokens are represented as $X_{v}=\{x_{1},x_{2},\dots,x_{n_{v}}\}$ and used as part of the LLM input alongside the text tokens $X_{q}=\{x_{n_{v}+1},x_{n_{v}+2},\dots,x_{n_{v}+n_{q}-1}\}$ that are embedded from tokenized text input by the embedding layer.
-Subsequently, the vision and text tokens are concatenated as the final input sequence and we denote it as $\{x_{i}\}_{1}^{t_{n}-1}$ where $t_{n}=n_{v}+n_{q}$, which is then fed into successive transformer layers.
-We denote the output of the $l$-th layer as $h_{t}^{(l)}$.
-Then, the vocabulary head $\zeta(\cdot)$ is used to predict the probability of the next token $x_{t}$ among the vocabulary set $\mathcal{X}$ as follows,
-$p(x_{t}\mid x_{<t})=\mathrm{softmax}(\zeta(h_{t}^{(L)}))_{x_{t}},x_{t}\in\mathcal{X}.$ (1)
-3.2 What Causes Hallucinations
-A hypothesis on the cause of hallucinations.
-Informed by the phenomenon of catastrophic forgetting (Zhai et al., 2023) in MLLMs, we argue that the capabilities of LLMs to comprehend and memorize different modalities are quite distinct.
-Taking image and text inputs as an example, since an image possesses a much higher information density than a piece of text, it is reasonable to assume that LLMs struggle to understand and memorize vision tokens compared to text tokens, prone to fantasies.
-Uncertainty quantification.
-Following the DoLa (Chuang et al., 2023), we compute the probability of the next token via the vocabulary head $\zeta$ on each layer during reasoning.
-Then, we introduce an entropy-based metric (Farquhar et al., 2024) to quantify the output uncertainty as $u=\sum-p_{k}\log p_{k}/\log{K}$, where $\{p_{k}\}_{i=1}^{K}$.
-With uncertainty, we make an important assumption.
-Assumption 3.1.
-LUFH: The Lower the Uncertainty, the Fewer the Hallucinations to be generated.
-We define $\gamma$ as the threshold that separates high uncertainty from low uncertainty in predictions.
-The candidate layer with uncertainty exceeding $\gamma$ is termed a premature layer.
-The proof of Assumption 3.1 (LUFH) is provided in Section 4.
-Based on the hypothesis, we aim to complement visual evidence in MLLMs to eliminate the hallucination caused by visual forgetting.
-In the following, we discuss our motivation, how MemVR is implemented, and why it can work in Section 3.3 and 3.4.
-3.3 Relationship Between Hallucinations and Uncertainty
-As findings of Chen et al.
-(2024) in LLMs: “incorrect tokens generally exhibit higher entropy than correct ones”, we also observe this phenomenon in MLLMs (visualization cases are shown in Appendix C.3).
-This implies the effectiveness of entropy-based metrics for detecting hallucinations.
-In this work, we use uncertainty as the metric.
-We present our in-depth findings in this section.
-Figure 3: Uncertainty of different early layers to predict the next token.
-Rows denote indices of the early layers, and column names are decoded tokens in each step.
-Uncertainty distribution is dynamic.
-Finding #1: In the context of tokens involving objects, attributes or relations, uncertainty is high.
-We conduct preliminary analysis with 32-layer LLaVA-1.5-7B.
-Specifically, we compute the uncertainty in the output distributions $p^{(l)}(\cdot\mid x_{<t})$ of early exiting layers.
-Figure 3 shows the uncertainty scores of different early layers when decoding the answer, we can observe that the computed uncertainty remains relatively high in later layers when predicting key entity objects, attributes, or relations, such as wooden, filled, and food in Figure 3.
-This phenomenon suggests that LLM is still uncertain about its predictions in the last few layers and may inject more factual knowledge into the predictions.
-On the other hand, when predicting function words and those tokens copied from the question, e.g., image, a, with, we observe that the uncertainty becomes very low from the middle layers.
-This finding implies that the model is deterministic for easy-to-predict tokens at the intermediate layer and keeps the distribution of outputs almost constant at higher layers, however, it is more uncertain for difficult-to-predict key tokens and may constantly change its predictions until the final layer.
-Finding #2: The uncertainty distribution of different tokens is dynamic during the generation.
-As can be seen in Figure 3, when visual factual knowledge is required for the prediction of the next token, such as wodden, filled, food, most layers of LLaVA are of higher uncertainty, and appear to shift the prediction of the later layers.
-The prediction of the next tokens such as quantity, target category, color, relation, etc.
-requires evident visual knowledge, but not for function words or commas
-From a qualitative perspective, intermediate values of uncertainty reflect the thinking or revision process of MLLMs, while too high uncertainty implies that the model is confused and relies on random guesses.
-For complex input images, the uncertainty of the output distributions tends to rise, which can propagate through subsequent layers, resulting in incorrect predictions, or gibberish.
-Figure 4: The illustration of how dynamic premature layer injection works.
-3.4 Visual Retracing and Its Dynamics
-Motivated by the above findings, we propose re-injecting visual evidence during elevated uncertainty in the model’s reasoning.
+Inspired by the common cognition: when the memory of an image seen the moment before is forgotten, people will look at it again for factual answers, we design look-twice mechanism, which treats visual tokens as supplementary evidences, re-injecting them into MLLMs through FFN as “key-value memory” at the middle trigger layer.
+This look-twice mechanism occurs when the model exhibits high uncertainty during inference, effectively enhancing factual alignment and modality balance.
+All details are listed below.
+4.1 Preliminary: Reformulation of FFN
+Vanilla FFN comprises two fully connected layers with non-linear activation in between. We suppose $\boldsymbol{x}\in\mathbb{R}^{d}$ as an input token of the FFN, and FFN function can be formulated as
+$\operatorname{FFN}(\boldsymbol{x})=\phi\left(\boldsymbol{x}\boldsymbol{W}_{1}\right)\boldsymbol{W}_{2}^{\top},$ (2)
+where $\phi$ is activation function like ReLU or SiLU (Liu et al., 2020), and $\boldsymbol{W}_{1},\boldsymbol{W}_{2}\in\mathbb{R}^{d\times D}$ are the weight matrices, in usual $D=4d$.
+Peculiarly, $\boldsymbol{W}_{1}$ and $\boldsymbol{W}_{2}$ can be rewritten as
+$\boldsymbol{W}_{1}=(\boldsymbol{k}_{1},\boldsymbol{k}_{2},\ldots,\boldsymbol{k}_{D}),\boldsymbol{W}_{2}=(\boldsymbol{v}_{1},\boldsymbol{v}_{2},\ldots,\boldsymbol{v}_{D}),$ (3)
+where $\boldsymbol{k}_{i},\boldsymbol{v}_{i}\in\mathbb{R}^{d}$ denote entries of key and value, respectively. As a result, the FFN can be reformulated as
+$\operatorname{FFN}(\boldsymbol{x})=\sum\phi\left(\left\langle\boldsymbol{x},\boldsymbol{k}_{i}\right\rangle\right)\cdot\boldsymbol{v}_{i}\;.$ (4)
+Thus, the FFN function can be construed as using input $\boldsymbol{x}$ as a query to measure similarity with keys, find matching values, and gather values by similarity, which works like a key-value memory storing the factual knowledge as found in previous studies (Geva et al., 2021; Jie et al., 2024).
+4.2 FFN with Visual Retracing
+Motivated by the findings in section §3.2 and §3.4, we propose Visual Retracing (VR), i.e., reinjecting visual evidence into the middle layer of the text decoder during elevated uncertainty during reasoning.
 This strategy treats visual tokens as anchors to recalibrate off-target predictions and reduces uncertainties in object, attribute, relationship tokens.
 Experimental results also demonstrate that our method reduces uncertainty and alleviates hallucinations as shown in Figure 8.
-We term this schema of re-injecting visual evidence as “visual retracing” that is to be elaborated further in Section 3.4.1.
-Further, we design a dynamic injection strategy, detailed in Section 3.4.2, ensuring timely visual evidence when generating uncertain visual-reliant tokens (Figure 4).
-3.4.1 FFN with Visual Retracing
-We introduce the implementation of our proposed memory-space visual retracing method.
-Previous study (Geva et al., 2021) has found that FFN acts as a key-value memory storing factual knowledge.
-Inspired by the fact that FFN executes analogous retrieval from its key-value memory, we consider “visual retracing” to serve as a simplified and efficient information re-retrieval process (Jie et al., 2024).
-Concretely, given an hidden token $x\in\mathbb{R}^{d}$ and vision features $z_{v}=(z_{v,1},\dots,z_{v,n_{v}})^{\top}\in\mathbb{R}^{n_{v}\times d^{\prime}}$, $\mathcal{K}_{\propto}\left({z}_{v,i}\right),\mathcal{V}_{\propto}\left({z}_{v,i}\right)\in\mathbb{R}^{d}$ are the key and value corresponding to ${z}_{v,i}$, visual retracing follows,
-${\operatorname{FFN}}_{\circlearrowright}^{(l)}({x}\propto z_{v})=(1-\alpha)\operatorname{FFN}^{(l)}(x)+\alpha\underline{\Lambda}_{\circlearrowright}^{(l)}({z}_{v}\mid{x}),$ (2)
-where $\alpha$ denotes the injection ratio of visual memory (proportional to image complexity), FFN comprises two FC layers with non-linear activation in between and can be formulated as $\operatorname{FFN}({x})=\phi\left({x}{W}_{1}\right){W}_{2}^{\top}$, $\phi$ is activation function like ReLU or SiLU (Liu et al., 2020).
-Separately, the weight matrices can be rewritten as: ${W}_{1}=\left({k}_{1},{k}_{2},\ldots,{k}_{m}\right),{W}_{2}=\left({v}_{1},{v}_{2},\ldots,{v}_{m}\right)^{\top}$, in which ${k}_{i}\in\mathbb{R}^{d}$ and ${v}_{i}\in\mathbb{R}^{d}$ are entries of key and value, respectively.
-Thus, FFN can be interpreted as using input ${x}$ as the query to calculate its similarity with keys to search for matching values.
-Analogously, we consider a simple and efficient retrieval process for visual retracing on $l$-th premature layer as,
-$\underline{\Lambda}_{\circlearrowright}^{(l)}({z}_{v}\mid{x})=\sum\nolimits_{i=1}^{n_{v}}\phi\left(\left\langle{x},\mathcal{K}_{\propto}\left({z}_{v,i}\right)\right\rangle\right)\cdot\mathcal{V}_{\propto}\left({z}_{v,i}\right).$ (3)
-From the perspective of FFN, we regard the $\left(\mathcal{K}_{\propto}\left({z}_{v,i}\right),\mathcal{V}_{\propto}\left({z}_{v,i}\right)\right)$ as new key-value entries to supplement vision-related evidence.
-$\mathcal{K}_{\propto}$ and $\mathcal{V}_{\propto}$ are key functions for aligning the dimension between vision features $z_{v}$ and the hidden token $x$.
-In MLLMs, $\mathcal{K}_{\propto}$ and $\mathcal{V}_{\propto}$ are regarded as visual projectors, e.g., MLPs or Q-Former (Li et al., 2023a), for aligning features of different modalities.
-3.4.2 Dynamic Premature Layer Injection
-To magnify the effectiveness of visual retracing, the optimal premature layer should ideally be the layer most uncertain about probable answers to visual questions.
+The reason we call this pattern of reinjecting visual evidence “visual retracing” is that the model finds and refreshes key visual memories based on the hidden states in this process.
+In particular, inspired by the fact that FFN executes analogous retrieval from its key-value memory, we consider VR to serve as a simplified and efficient information re-retrieval process.
+Given a hidden token $\boldsymbol{x}\in\mathbb{R}^{d}$ and dimension-aligned vision tokens $\boldsymbol{z}_{v}$, FFN with visual retracing at $l$-th layer can be written as follows
+${\operatorname{FFN}}^{(l)}(\boldsymbol{x}\propto\boldsymbol{z}_{v})=\alpha\underline{\Delta}+(1-\alpha)\operatorname{FFN}^{(l)}(\boldsymbol{x}),$ (5)
+where $\boldsymbol{z}_{v}=(\boldsymbol{z}_{v,1},\dots,\boldsymbol{z}_{v,N_{v}})\in\mathbb{R}^{d\times N_{v}}$, ${x}\propto\boldsymbol{z}_{v}$ denotes execute VR $\underline{\Delta}$ from $\boldsymbol{x}$ to visual features $\boldsymbol{z}_{v}$, and $\alpha\in[0,1]$ denotes injection ratio of visual memory through the FFN layer which proportional to image complexity.
+Specifically, instead of performing retrieval via cross-attention layers as in previous approaches (Li et al., 2022; Alayrac et al., 2022), we consider a simple retrieval process for VR as,
+$\underline{\Delta}({\boldsymbol{z}}_{v}\mid\boldsymbol{x})=\sum\nolimits_{i=1}^{N_{v}}\phi(\left\langle\boldsymbol{x},{\boldsymbol{z}}_{v,i}\right\rangle)\cdot{\boldsymbol{z}}_{v,i}.$ (6)
+From the perspective of FFN, visual retracing works by treating $\boldsymbol{x}$ as a query, and $\left\langle{\boldsymbol{z}}_{v,i}:{\boldsymbol{z}}_{v,i}\right\rangle$ as new key-value entries (visual evidence) to supplement vision-related information in the hidden states.
+In this information re-retrieval process, MemVR does not introduce any parameters that need to be trained.
+Notably, since the size of key-value memory $D$ in FFN typically far exceeds the number of visual tokens $N_{v}$ (for instance, $D=11008$ in LLaMA-7B and $N_{v}=256$ for ViT-L/14, $N_{v}\ll D$), the computation of MemVR is negligible.
+Thus, VR operation is more efficient than the cross-attention mechanism with quadratic complexity.
+4.3 Dynamic Triggered Visual Retracing
+To magnify the effectiveness of VR, the optimal strategy should be to trigger VR dynamically based on the token, and the selection of the trigger layer should also be dynamically determined.
 In practice, we consider that the uncertainty of a candidate layer exceeding the threshold $\gamma$ warrants visual retracing.
-Inspired by the fact that early exit patterns (Teerapittayanon et al., 2016; Elbayad et al., 2020; Schuster et al., 2022) have proven effective in directly employing the language heads $\zeta$ to the hidden states of the middle layers, even without a special training process (Kao et al., 2020), we compute the uncertainty of the next token probability on the early layers for reasoning.
-As our Finding #2, we utilize layer-specific uncertainty to allow for dynamic premature layer injection at each time step as illustrated in Figure 3.
-Algorithm 1 Dynamic Injection Strategy
-1:$h_{t},{z}_{v}$ denote hidden states and visual evidence. We set trigger = True.
-2:for $l$ = 1 to $L-1$ do
-3:     $p^{(l)}=\mathrm{softmax}(\zeta(h_{t}^{(l)}))_{x_{t}}$
-4:     $u^{(l)}={\sum-p^{(l)}\log p^{(l)}}/{\log{K}}.$
+Inspired by the fact that early exit patterns (Elbayad et al., 2020; Schuster et al., 2022) have proven effective in directly employing the language heads $\zeta$ to the hidden states of the middle layers, even without a special training process (Kao et al., 2020), we compute the uncertainty of the next token probability on the early layers for reasoning.
+We utilize layer-specific uncertainty to allow for dynamic premature layer injection at each time step.
+Dynamic Triggered MemVR.
+For MLLMs with different numbers of layers, as Algorithm 1 shown, the dynamic triggered MemVR strategy called MemVR-dynamic, identifies desirable premature layers among the candidate layers for visual retracing based on output uncertainty of different layers, thus better amplifying the effect of visual retracing.
+Algorithm 1 Dynamic Triggered MemVR Strategy
+0:  MLLM $\mathcal{M}_{\theta}^{\text{MLLM}}$, text query $x$, image input $v$.
+0:  Model response $y_{t}^{b}$.
+1:  At every decoding step $t$:
+2:  Initial set to trigger = True.
+3:  for $l$ = 1 to $L-1$ do
+4:     $u^{(l)}={\sum-p_{\theta}^{(l)}\log p_{\theta}^{(l)}}/{\log{N}}.$   % uncertainty
 5:     if trigger = = True and $u^{(l)}>\gamma$ then
-6:         Execute $\underline{\Lambda}_{\circlearrowright}^{(l+1)}({z}_{v}\mid{h_{t}^{(l+1)}})$
-7:         Select ${\operatorname{FFN}}_{\circlearrowright}^{(l+1)}({h_{t}^{(l+1)}}\propto z_{v})$
-8:         trigger=False # only once
-9:     else
-10:         Select ${\operatorname{FFN}}^{(l+1)}({h_{t}^{(l+1)}})$
-11:     end if
-12:end for
-For MLLMs with different numbers of layers, we partition the layers into several (typically three) buckets according to the total layers for sensible visual retracing triggers, as detailed in Appendix C.1.
-This dynamic injection strategy identifies desirable premature layers for visual retracing based on output uncertainty, thus better leveraging insights from different layers.
-Static Fixed Layer Injection.
-In addition to the dynamic premature layer injection strategy, another more straightforward strategy worth considering is to perform a brute-force experiment on all possible early layers using a validation set and selecting the layer with the best average performance.
+6:        Execute $\underline{\Delta}({\boldsymbol{z}}_{v}\mid{h_{t}^{(l+1)}})$4, Eq. (6)}
+7:        Select ${\operatorname{FFN}}^{(l+1)}({h_{t}^{(l+1)}}\propto\boldsymbol{z}_{v})$4, Eq. (5)}
+8:        trigger = False   % set to only look-twice
+9:     end if
+10:  end for
+11:  $\mathcal{M}_{\theta}^{\text{MLLM}}$ decoding, obtain current token $y_{t}^{b}$.
+Static Triggered MemVR.
+Besides MemVR-dynamic, another more straightforward strategy worth considering is to perform a brute-force experiment on all possible early layers using a validation set and selecting the layer with the best average performance.
 We refer to this simple strategy as MemVR-static.
-However, the MemVR-static approach presents two limitations: (I) it requires more extensive hyperparameter tuning across layers, and (II) the optimal layer is highly sensitive to data distribution, necessitating an in-distribution validation set.
-In contrast, our proposed dynamic layer injection strategy mitigates these challenges by reducing the layer search space and improving robustness without depending on in-distribution validation.
-Empirical comparisons between our MemVR using dynamic and static strategies are provided in Section 4.2 and Table 4.
-3.5 Theoretically Understanding Why MemVR Works
-In order to gain further insight into the reasons behind the effectiveness of MemVR in mitigating hallucinations and its robust performance on general benchmarks, we attempt to explain these phenomena via the following three theorems from an information-theoretic perspective.
+Nonetheless, MemVR-static exhibits two notable limitations.
+Firstly, it demands more comprehensive hyperparameter tuning across different layers.
+Secondly, the optimal layer is extremely sensitive to the data distribution, which mandates the use of an in-distribution validation set.
+MemVR-dynamic vs.
+-static.
+In contrast to MemVR-static, MemVR-dynamic strategy effectively alleviates these challenges.
+It achieves this by narrowing down the layer search space and enhancing robustness, all without relying on an in-distribution validation set.
+Empirical comparisons between the proposed MemVR implemented with the dynamic and static strategies are presented in Section §5.4 and Table 10.
+4.4 Theoretical Analysis
+To further understand why MemVR effectively mitigates hallucinations and performs robustly on general benchmarks, we explain these phenomena using three theorems below.
 Theorem 4.1.
-Let ${H}_{vq}$ be the hidden states of FFN and $\hat{H}_{vq}$ be after reinjection of visual evidence $Z_{v}$.
-MemVR enhances Mutual Information (MI) between $\hat{H}_{vq}$ and $Z_{v}$: $I(\hat{H}_{vq};Z_{v})\geq I(H_{vq};Z_{v})$.
-The reinjection of $Z_{v}$ at intermediate layers of the model facilitates the replenishment of critical visual information, which may have been lost or distorted through earlier layers.
-This process increases MI between the hidden states and the visual features, ensuring adequate visual context is preserved.
+Let $\boldsymbol{x}$ be the hidden states of FFN and $\hat{\boldsymbol{x}}$ be after reinjecting visual evidence $\boldsymbol{z}_{v}$.
+MemVR enhances Mutual Information (MI) between $\hat{\boldsymbol{x}}$ and $\boldsymbol{z}_{v}$ as:
+$I(\hat{\boldsymbol{x}};\boldsymbol{z}_{v})\geq I(\boldsymbol{x};\boldsymbol{z}_{v}).$ (7)
 Theorem 4.2.
-Let $Y$ be the target output dependent on hidden states.
-If MI between ${H}_{vq}$ and ${Z}_{v}$ increases, then conditional entropy $H(Y\mid{H}^{(l)}_{vq})$ decreases with $H(Y\mid\hat{H}_{vq})\leq H(Y\mid H_{vq})$.
-According to Theorem 3.1, 3.2, and DPI (Cover et al., 1991), MemVR improves the quality of the representations $H_{vq}$ and reduces the uncertainty in the output $Y$.
-As a result, the probability of hallucinations decreases, which is consistent with the observed findings of hallucination mitigation.
+Let $\boldsymbol{y}$ be the target output dependent on hidden states.
+If MI between $\boldsymbol{x}$ and $\boldsymbol{z}_{v}$ increases, then conditional entropy $H(\boldsymbol{y}\mid\boldsymbol{x})$ decreases with
+$H(\boldsymbol{y}\mid\hat{\boldsymbol{x}})\leq H(\boldsymbol{y}\mid\boldsymbol{x}).$ (8)
 Theorem 4.3.
-Within the Information Bottleneck (IB) framework, the loss of objective function, represented by the notation $\mathcal{L}(T)$, is optimized by MemVR, which is defined as $\mathcal{L}(\hat{H}_{vq})\leq\mathcal{L}(H_{vq})$, where $\mathcal{L}(H_{vq})=I(H_{vq};X_{vq})-\beta I(H_{vq};Y)$ is IB loss, and $\ beta$ is a trade-off parameter.
-Anchored in the information bottleneck framework, MemVR optimizes the delicate balance between retaining relevant information from multimodal inputs and compressing non-essential details, thereby safeguarding the predictive performance of the hidden representations for the target output $Y$.
-The theoretical underpinnings of MemVR are supported by the Data Processing Inequality (Cover et al., 1991) and the contraction properties of stochastic mappings in deep neural networks, as shown in various studies on the Information Bottleneck Principle (Achille & Soatto, 2018).
-By enhancing mutual information and reducing the uncertainty in hidden states, MemVR effectively mitigates hallucinations while preserving computational efficiency.
-Detailed proofs are in Appendix B.
-Method   MSCOCO   A-OKVQA   GQA
-%Accuracy %F1 Score %Accuracy %F1 Score %Accuracy %F1 Score
-LLaVA1.5-7B 81.38  $\uparrow$0.0 79.65  $\uparrow$0.0 79.13  $\uparrow$0.0 79.10  $\uparrow$0.0 79.00  $\uparrow$0.0 79.13  $\uparrow$0.0
-+ VCD (Leng et al., 2024) 84.66  $\uparrow$3.3 84.52  $\uparrow$4.9 80.99  $\uparrow$1.8 82.30  $\uparrow$3.2 81.74  $\uparrow$2.7 82.16  $\uparrow$3.0
-+ OPERA (Huang et al., 2024a) 84.77  $\uparrow$3.4 85.46  $\uparrow$5.8 84.27  $\uparrow$5.1 84.08  $\uparrow$5.0 84.03  $\uparrow$5.0 83.83  $\uparrow$4.7
-+ MemVR (Ours) 87.00  $\uparrow$5.7 85.87  $\uparrow$6.2 86.21  $\uparrow$7.0 86.64  $\uparrow$7.5 85.25  $\uparrow$6.2 85.59  $\uparrow$6.4
-Qwen-VL-10B 83.79  $\uparrow$0.0 81.13  $\uparrow$0.0 84.74  $\uparrow$0.0 83.27  $\uparrow$0.0 84.41  $\uparrow$0.0 82.66  $\uparrow$0.0
-+ VCD (Leng et al., 2024) 84.27  $\uparrow$0.4 82.12  $\uparrow$1.0 84.09  $\downarrow$0.7 82.53  $\downarrow$0.7 83.73  $\downarrow$0.7 82.75  $\uparrow$0.1
-+ OPERA (Huang et al., 2024a) 84.93  $\uparrow$1.1 83.41  $\uparrow$2.3 - - - -
-+ MemVR (Ours) 84.07  $\uparrow$0.3 81.55  $\uparrow$0.4 86.43  $\uparrow$1.8 85.56  $\uparrow$2.3 85.69  $\uparrow$1.3 84.53  $\uparrow$1.9
-Table 2: Performance evaluation on POPE.
-The best results in each scenario are bolded for clarity.
-We report the averages under the three settings, e.g., Random, Popular, and Adversarial to show the robustness of the different methods directly.
-Green denotes improvement, and Red means degradation.
-Method Commonsense QA(Reasoning)   Object-level Hallucination   Attribute-level Hallucination  Total Scores
-Existence Count Position Color
-LLaVA1.5-7B 110.71  $\uparrow$0.0 175.67  $\uparrow$0.0 124.67  $\uparrow$0.0 114.00  $\uparrow$0.0 151.00  $\uparrow$0.0 676.05
-+ VCD (Leng et al., 2024) 112.86  $\uparrow$9.9 184.66  $\uparrow$9.0 138.33  $\uparrow$13.6 128.67  $\uparrow$14.6 153.00  $\uparrow$2.0 717.52
-+ OPERA (Huang et al., 2024a) 115.71  $\uparrow$5.5 180.67  $\uparrow$5.0 133.33  $\uparrow$8.6 123.33  $\uparrow$9.3 155.00  $\uparrow$4.0 708.04
-+ MemVR (Ours) 121.42  $\uparrow$18.5 190.00  $\uparrow$14.3 155.00  $\uparrow$30.3 133.33  $\uparrow$19.3 170.00  $\uparrow$19.0 769.75
-Qwen-VL-10B 106.40  $\uparrow$0.0 155.00  $\uparrow$0.0 127.67  $\uparrow$0.0 131.67  $\uparrow$0.0 173.00  $\uparrow$0.0 693.74
-+ VCD (Leng et al., 2024) 104.33  $\downarrow$2.1 156.00  $\uparrow$1.0 131.00  $\uparrow$3.3 128.00  $\downarrow$3.6 181.67  $\uparrow$8.6 701.00
-+ OPERA (Huang et al., 2024a) 104.33  $\uparrow$2.2 165.00  $\uparrow$6.9 145.00  $\uparrow$4.8 133.33  $\uparrow$1.6 180.00  $\uparrow$7.0 727.66
-+ MemVR (Ours) 120.00  $\uparrow$13.6 185.00  $\uparrow$30.0 145.00  $\uparrow$17.3 123.33  $\downarrow$8.3 185.00  $\uparrow$12.0 758.33
-Table 3: Results on the hallucination subset of MME (including commonsense reasoning, existence, count, position, color scores). The best are in bold. MemVR achieves dramatic improvements.
+Within the Information Bottleneck (IB) framework, the loss of objective function, represented by the notation $\mathcal{L}(T)$, is optimized by MemVR, which is defined as $\mathcal{L}(\hat{\boldsymbol{x}})\leq\mathcal{L}(\boldsymbol{x})$, where $\mathcal{L}(\boldsymbol{x})=I(\boldsymbol{x};\boldsymbol{c})-\beta I(\boldsymbol{x};\boldsymbol{y})$ is IB loss, $\boldsymbol{c}$ denotes input embedding and $\beta$ is a trade-off parameter.
+Intuition: The proofs are provided in Appendix A.
+The theoretical basis of MemVR draws support from the DPI (Cover et al., 1991) and the contraction properties of stochastic mappings in deep neural networks, as evidenced in numerous IB-related studies (Achille & Soatto, 2018).
+By enhancing MI and reducing uncertainty in hidden states, MemVR effectively alleviates hallucination while maintaining efficiency.
 5 Experiments
-This section details the evaluation of our MemVR across three MLLMs on seven benchmarks.
 5.1 Experiment Setup
-Datasets and Metrics.
-To rigorously assess the effectiveness of our proposed method, we conduct a comprehensive set of experiments across two benchmarks specifically designed to evaluate hallucination mitigation and five general-purpose benchmarks to gauge the general performance:
-  • Hallucination benchmarks: Polling-based Object Probing Evaluation (POPE) (Li et al., 2023c), and Caption Hallucination Assessment with Image Relevance (CHAIR) (Rohrbach et al., 2018);
-  • General-purpose benchmarks: VizWiz-VQA (Gurari et al., 2018), MLLM Comprehensive Evaluation (MME) (Fu et al., 2023), Multimodal Benchmark (MMBench) (Liu et al., 2023b), Multimodal Veterinarian (MM-Vet) (Yu et al., 2024c), LLaVA-Bench (in-the-wild) (Liu et al., 2024).
-More detailed information on these various benchmarks can be obtained from the Appendix C.1.
-Backbones and Baselines.
-To evaluate our method, we utilize three well-known MLLMs: LLaVA-1.5 (Liu et al., 2024), Qwen-VL (Bai et al., 2023), and GLM4V (Wang et al., 2023).
-Further, We compare our methods with classic training-tree SOTA methods designed to mitigate object hallucination, including visual contrastive decoding SOTA VCD Leng et al.
-(2024), OPERA (Huang et al., 2024a) based on overconfidence penalty and hindsight allocation.
-As Dola (Chuang et al., 2023) is layer-wise contrastive decoding for LLMs and performs poorly in MLLMs, it will not be shown in the experiment.
-Experimental results are obtained and benchmarked using unified implementation.
-Implementation Details.
-Greedy search is used as the default decoding strategy in MemVR for all benchmarks.
-For benchmarks, annotation questions are adapted to MLLM templates.
-For POPE, COCO, A-OKVQA, and GQA are used, while MMBench_DEV_EN is used for MMBench.
-MM-Vet is assessed using MM-Vet Online Evaluator, and gpt4-1106-preview is used for LLaVA-Bench.
-CHAIR uses images from COCO Val2014 with the query ”Please describe this image in detail”.
-In MemVR, do_sample=False, temperature=0, threshold=0.75, beam=1.
-All settings of the compared method follow the default configurations from the original papers.
-More details are in Appendix C.1.
-Figure 5: Results on MMBench. MemVR enhances comprehensive performance on diverse tasks.
-Strategy Static-3 Static-7 Static-11 Static-15 Static-19 Static-23 Static-27 Static-31 Dynamic
-Preception 1508.08 1500.48 1526.40 1529.07 1513.96 1500.29 1510.47 1510.34 1512.80
-Cognition 364.64 347.14 362.86 352.14 350.36 357.86 355.71 346.43 383.92
-Total Score 1872.72 1847.62 1889.26 1881.22 1864.32 1858.15 1866.18 1856.77 1896.72
-Table 4: Results of MemVR-static and MemVR-dynamic on MME. Static-# indicates fixation on layer # for visual retracing. MemVR-dynamic achieves optimal performance improvements.
-4.2 Quantitative Results
-Key Finding: MemVR consistently outperforms the baselines in mitigating hallucinations and improving overall accuracy across various scenarios.
-Figure 6: Comparison of different injection ratios $\alpha$ on cognition scores.
-Q1: How does MemVR performance on hallucination benchmark?
-We conduct POPE, CHAIR, and MME evaluations, as shown in Table 2, Table 3 and Table 5, our MemVR obviously surpasses all of the compared baselines.
-For the results of POPE evaluation, we observe that our proposed MemVR presents robust effects.
-The performance of MemVR surpasses the baseline results by large margins, i.e., average up to +7.0% accuracy and +7.5% F1 score on A-OKVQA dataset under Random, Popular, and Adversarial settings.
-As showcased in Table 5, our MemVR achieves up to 15.6% improvements on CHAIRI compared with vanilla LLaVA-1.5-7B.
-For MME subset evaluations (encompass both object-level and attribute-level hallucinations), results in Table 3 show that MemVR achieves a uniform improvement in handling object-level and attribute-level hallucinations, as well as commonsense reasoning.
-Existence, Count, and Color scores all achieve dramatic improvements (Existence score in Qwen-VL up 30).
-On the contrary, Position scores are relatively low, which suggests weak position reasoning capability in MLLMs.
-Model Length CHAIRS $\downarrow$ CHAIRI $\downarrow$ Recall $\uparrow$
-LLaVA-1.5 100.6 50.0  $\uparrow$0.0 15.4  $\uparrow$0.0 77.1  $\uparrow$0.0
-+ VCD 100.4 48.6  $\downarrow$1.4 14.9  $\downarrow$0.5 77.3  $\uparrow$0.2
-+ OPERA 98.6 47.8  $\downarrow$2.2 14.6  $\downarrow$0.8 76.8  $\downarrow$0.3
-+ MemVR 99.6 46.6  $\downarrow$3.4 13.0  $\downarrow$2.4 80.8  $\uparrow$3.7
-Table 5: CHAIR hallucination evaluation results of LLaVA. Small values correspond to fewer hallucinations.
-Q2: How well does MemVR perform on general-purpose benchmarks?
-We evaluate the performance of MemVR on general-purpose benchmarks, i.e., VizWiz, MME, MMBench, MM-Vet, and LLaVA-Bench.
-Appx.
-C summarizes the results on MMBench, highlighting MemVR’s comparative performance relative to SOTA methods.
-As shown in Table 6, MemVR consistently outperforms competing models.
+Datasets and Evaluation.
+To rigorously assess the effectiveness of our proposed MemVR, we conduct a comprehensive set of experiments across POPE benchmark (Li et al., 2023b), CHAIR (Rohrbach et al., 2018), VizWiz-VQA (Gurari et al., 2018), MME (Fu et al., 2023), MMBench (Liu et al., 2023b), MM-Vet (Yu et al., 2024b), LLaVA-Bench (in-the-wild) (Liu et al., 2024b), HallusionBench (Guan et al., 2024).
+More details can be found in Appendix B.1.
+Implementation Details. Usually, we set $\gamma$=0.75. All settings of baseline methods follow the default configurations from the original papers. More details are in Appendix B.2.
+Table 3: CHAIR evaluation results of different methods.
+Methods CHAIRS $\downarrow$ CHAIRI $\downarrow$ Average $\downarrow$ Len Recall $\uparrow$
+LLaVA-1.5 50.0  $\downarrow$0.0 15.4  $\downarrow$0.0 32.7  $\downarrow$0.0 100.6 77.1  $\uparrow$0.0
++ OPERA 47.8  $\downarrow$2.2 14.6  $\downarrow$0.8 31.2  $\downarrow$0.5 98.6 76.8  $\downarrow$0.3
++ VCD 48.6  $\downarrow$1.4 14.9  $\downarrow$0.5 31.8  $\downarrow$0.5 100.4 77.3  $\uparrow$0.2
++ ICD 56.2  $\uparrow$6.2 16.3  $\uparrow$0.9 36.3  $\uparrow$3.6 103.4 16.3  $\downarrow$60.
++ MemVR 46.6  $\downarrow$3.4 13.0  $\downarrow$2.4 29.8  $\downarrow$0.5 99.6 80.8  $\uparrow$3.7
+Table 4: Performance evaluation on POPE benchmark.
+The best results are in green.
+We report accuracy and f1-score under three settings, e.g., Random, Popular, Adversarial, and Average, to show the robustness of the different methods directly.
+Evaluation Methods  Random  Popular  Adversarial  Average
+Accuracy $\uparrow$ F1-score $\uparrow$ Accuracy $\uparrow$ F1-score $\uparrow$ Accuracy $\uparrow$ F1-score $\uparrow$ Accuracy $\uparrow$ F1-score $\uparrow$
+MSCOCO LLaVA-1.5-7B 83.49  $\uparrow$0.0 82.28  $\uparrow$0.0 79.98  $\uparrow$0.0 79.34  $\uparrow$0.0 76.03  $\uparrow$0.0 76.26  $\uparrow$0.0 79.83  $\uparrow$0.0 79.29  $\uparrow$0.0
+OPERA (Huang et al., 2024a) 87.53  $\uparrow$4.0 86.45  $\uparrow$4.2 84.21  $\uparrow$4.2 83.50  $\uparrow$4.2 80.88  $\uparrow$4.9 80.69  $\uparrow$4.4 84.21  $\uparrow$4.4 83.55  $\uparrow$4.3
+ICD (Wang et al., 2024) 84.87  $\uparrow$1.4 83.27  $\uparrow$1.0 82.93  $\uparrow$3.0 81.45  $\uparrow$2.1 81.07  $\uparrow$5.0 79.96  $\uparrow$3.7 82.96  $\uparrow$3.1 81.56  $\uparrow$2.3
+VCD (Leng et al., 2024) 86.84  $\uparrow$3.4 86.83  $\uparrow$4.6 82.65  $\uparrow$2.7 83.37  $\uparrow$4.0 77.31  $\uparrow$1.3 79.28  $\uparrow$3.0 82.27  $\uparrow$2.4 83.16  $\uparrow$3.9
+MemVR (ours) 88.50  $\uparrow$5.0 87.34  $\uparrow$5.0 87.10  $\uparrow$7.1 86.01  $\uparrow$6.7 85.20  $\uparrow$9.2 84.28  $\uparrow$8.0 86.93  $\uparrow$7.1 85.88  $\uparrow$6.6
+A-OKVQA LLaVA-1.5-7B 83.45  $\uparrow$0.0 82.56  $\uparrow$0.0 79.90  $\uparrow$0.0 79.59  $\uparrow$0.0 74.04  $\uparrow$0.0 75.15  $\uparrow$0.0 79.13  $\uparrow$0.0 79.10  $\uparrow$0.0
+OPERA (Huang et al., 2024a) 88.27  $\uparrow$4.8 87.54  $\uparrow$5.0 85.17  $\uparrow$5.3 84.74  $\uparrow$5.2 79.37  $\uparrow$5.3 79.97  $\uparrow$4.8 84.27  $\uparrow$5.1 84.08  $\uparrow$5.0
+ICD (Wang et al., 2024) 85.57  $\uparrow$2.1 85.06  $\uparrow$2.5 81.93  $\uparrow$2.0 81.95  $\uparrow$2.4 77.43  $\uparrow$3.4 78.99  $\uparrow$3.8 81.64  $\uparrow$2.5 82.00  $\uparrow$2.9
+VCD (Leng et al., 2024) 86.15  $\uparrow$2.7 86.34  $\uparrow$3.8 81.85  $\uparrow$2.0 82.82  $\uparrow$3.2 74.97  $\uparrow$0.9 77.73  $\uparrow$2.6 80.99  $\uparrow$1.9 82.30  $\uparrow$3.2
+MemVR (ours) 91.10  $\uparrow$7.7 90.83  $\uparrow$8.3 87.33  $\uparrow$7.4 87.43  $\uparrow$7.8 80.20  $\uparrow$6.2 81.66  $\uparrow$6.5 86.21  $\uparrow$7.1 86.64  $\uparrow$7.5
+GQA LLaVA-1.5-7B 83.73  $\uparrow$0.0 82.95  $\uparrow$0.0 78.17  $\uparrow$0.0 78.37  $\uparrow$0.0 75.08  $\uparrow$0.0 76.06  $\uparrow$0.0 78.99  $\uparrow$0.0 79.13  $\uparrow$0.0
+OPERA (Huang et al., 2024a) 88.27  $\uparrow$4.5 87.52  $\uparrow$4.6 83.07  $\uparrow$4.9 82.93  $\uparrow$4.6 80.77  $\uparrow$5.7 81.05  $\uparrow$5.0 84.04  $\uparrow$5.1 83.83  $\uparrow$4.7
+ICD (Wang et al., 2024) 84.90  $\uparrow$1.2 84.22  $\uparrow$1.3 78.37  $\uparrow$0.2 78.81  $\uparrow$0.4 75.97  $\uparrow$0.9 76.93  $\uparrow$0.9 79.75  $\uparrow$0.8 79.99  $\uparrow$0.9
+VCD (Leng et al., 2024) 86.65  $\uparrow$2.9 86.99  $\uparrow$4.0 80.73  $\uparrow$2.6 82.24  $\uparrow$3.9 76.09  $\uparrow$1.0 78.78  $\uparrow$2.7 81.16  $\uparrow$2.2 82.67  $\uparrow$3.5
+MemVR (ours) 89.60  $\uparrow$5.9 89.32  $\uparrow$6.4 84.63  $\uparrow$6.5 84.98  $\uparrow$6.6 81.53  $\uparrow$6.4 82.48  $\uparrow$6.4 85.25  $\uparrow$6.3 85.59  $\uparrow$6.5
+Table 5: HallusionBench evaluation results of different methods.
+Methods fACC $\uparrow$ qACC $\uparrow$ ${easy}$A $\uparrow$ ${hard}$A $\uparrow$ aACC $\uparrow$
+LLaVA-1.5 17.9  $\uparrow$0.0 8.13  $\uparrow$0.0 36.0  $\uparrow$0.0 36.7  $\uparrow$0.0 41.5  $\uparrow$0.0
++ OPERA 16.2  $\downarrow$1.8 5.49  $\downarrow$2.6 37.6  $\uparrow$0.9 35.4  $\downarrow$1.3 41.2  $\downarrow$0.3
++ ICD 13.9  $\downarrow$4.0 8.35  $\uparrow$0.2 36.9  $\uparrow$0.2 33.5  $\downarrow$3.4 38.2  $\downarrow$3.3
++ VCD 13.9  $\downarrow$4.0 11.4  $\uparrow$3.3 33.0  $\downarrow$3.7 34.7  $\downarrow$3.0 41.1  $\downarrow$0.4
++ MemVR 17.9  $\uparrow$0.0 9.01  $\uparrow$0.9 36.9  $\uparrow$0.9 37.7  $\uparrow$1.0 42.5  $\uparrow$1.0
+5.2 Results on Hallucination Benchmarks
+We conduct hallucination evaluations on CHAIR, POPE, and HallusionBench with results presented in Table 3, Table 4 and Table 5.
+In the POPE evaluation, MemVR demonstrates robust effects.
+Its performance significantly exceeds baseline results, with an average accuracy increase of up to +7.0% and an F1-score increase of up to +7.5% on the A-OKVQA dataset under the Random, Popular, and Adversarial settings.
+MemVR clearly outperforms all compared SOTA methods.
+As shown in Table 3, compared with vanilla LLaVA-1.5, MemVR achieves 6.8% and 15.6% improvement on CHAIRS and CHAIRI metrics.
+Table 5 shows the HallusionBench evaluation results, where MemVR outperforms vanilla LLaVA-1.5 and other compared methods, achieving the best performance and consistent improvement across different metrics such as $hard$aACC and aACC.
+5.3 Results on General-purpose Benchmarks
+We evaluate the performance of MemVR on general-purpose benchmarks, i.e., LLaVA-Bench, MM-Vet, MME, MMBench, and VizWiz.
+As shown in Table 6, MemVR consistently outperforms competing models on LLaVA-Bench.
 Besides, MemVR achieves a significant improvement in overall performance listed in Table 7, with an average increase of 6.1% in OCR and spatial awareness tasks, demonstrating superior generalization capabilities.
-These results indicate that compared with CD-based methods, MemVR excels in hallucination mitigation and delivers competitive performance on general-purpose benchmarks.
-More complete results are in Appendix C.
-Model Convs $\uparrow$ Detail $\uparrow$ Complex $\uparrow$ All $\uparrow$
-LLaVA-1.5 58.8  $\uparrow$0.0 52.1  $\uparrow$0.0 74.6  $\uparrow$0.0 63.4  $\uparrow$0.0
-+ VCD 57.8  $\downarrow$1.0 50.8  $\downarrow$1.3 77.9  $\uparrow$3.3 59.1  $\downarrow$4.3
-+ OPERA 59.5  $\uparrow$0.7 49.6  $\downarrow$2.5 78.6  $\uparrow$4.0 59.8  $\downarrow$3.6
-+ MemVR 63.8  $\uparrow$0.5 52.6  $\uparrow$0.5 77.9  $\uparrow$3.3 64.0  $\uparrow$0.6
-Table 6: LLaVA-Bench evaluation results.
-Q3: How efficient is MemVR?
-MemVR operates dynamically based on the uncertainty, which employs visual retracing when the uncertainty exceeds threshold $\gamma$ on the early layer.
-If the uncertainty remains low across all layers—indicating that the model is highly confident in its generated results—MemVR is not triggered.
-This mechanism ensures efficient inference without extra computational overhead.
-Compared with VCD and OPERA, they need inference twice or the rollback strategy leads to exponentially added overheads, our MemVR only once regular inference.
-Figure 7: MME evaluation results under different mixing ratios of noise and visual features.
-Q4: How does visual uncertainty affect MemVR?
-To examine the impact of visual uncertainty on MemVR ’s performance, we introduced Gaussian noise (McHutchon & Rasmussen, 2011) into the extracted visual features.
-We gradually increased the noise level to assess how MemVR ’s performance would respond, using MME as the benchmark and LLaVA as the backbone model.
-As illustrated in Figure 7, both the perception and cognition scores declined as the noise step increased.
-At high noise, the performance fell significantly below the baseline.
-This demonstrates that MemVR is sensitive to the quality of visual features and is efficient at understanding shallow information.
-Figure 8: Results of GLM4V-9B. MemVR enhances comprehensive performance on diverse tasks.
-Method gn_kw_rec rec ocr_sp ocr ocr_sp_rec ocr_kw_rec ocr_gn_sp Total
-LLaVA1.5-7B 18.1  $\uparrow$0.0 67.6  $\uparrow$0.0 17.7  $\uparrow$0.0 48.3  $\uparrow$0.0 60.0  $\uparrow$0.0 21.2  $\uparrow$0.0 10.0  $\uparrow$0.0 31.1  $\uparrow$0.0
-+ VCD 19.2  $\uparrow$1.1 62.2  $\downarrow$5.4 15.8  $\downarrow$1.9 29.2  $\downarrow$20.0 42.5  $\downarrow$17.5 17.5  $\downarrow$3.7 60.0  $\uparrow$50.0 30.2  $\downarrow$1.1
-+ OPERA 21.8  $\uparrow$3.7 61.9  $\downarrow$5.7 21.5  $\uparrow$3.8 51.7  $\uparrow$3.4 56.2  $\downarrow$3.8 11.2  $\downarrow$10.0 30.0  $\uparrow$1.4 32.0  $\uparrow$0.9
-+ MemVR 19.5  $\uparrow$1.4 70.3  $\uparrow$2.7 23.8  $\uparrow$6.1 48.3  $\uparrow$0.0 58.8  $\downarrow$1.2 21.2  $\uparrow$0.0 30.0  $\uparrow$20.0 32.4  $\uparrow$1.3
-Table 7: MM-Vet evaluation results with multiple complicated multimodal tasks, where gn denotes language generation, kw means knowledge, sp denotes spatial awareness, and rec is recognition.
-Figure 9: Visualisation of uncertainty across layers without and with MemVR. MemVR effectively reduces uncertainty after 8th layer, contributing to hallucination mitigations as Assumption 3.1.
-Q5: How does layer uncertainty assist with visual retracing?
-MemVR leverages layer entropy to trigger visual retracing dynamically.
-We compared the performance of fixed retracing layers with dynamic retracing layers and found that dynamic retracing outperformed fixed retracing, as demonstrated in Table 4.
-Under all experimental conditions, dynamic retracing achieved the highest total score on the MME evaluation using LLaVA.
-Furthermore, we analyzed the injection ratio for MemVR, with the optimal injection ratio observed around 0.15, as shown in Table 4.
-This indicates that a balanced level of $\alpha$ is crucial for maximizing performance.
-Q6: Can MemVR adapt to other MLLMs?
-Definitely, MemVR is designed to be flexible and compatible with various architectures.
-A key advantage of MemVR lies in its ability to function without requiring modifications to the Transformers library, facilitating smooth integration into both older and cutting-edge models.
-We conducted extensive experiments across multiple benchmarks with LLaVA, Qwen-VL, and GLM-4V, and all three models outperformed baselines.
-4.3 Qualitative Study: What types of faults can our method address?
-Figure 10: A case study in long text generation. MemVR effectively mitigates hallucinations.
-In addition to evaluating single-word question-answering (QA) benchmarks, we further explore the models’ ability to generate comprehensive long-text descriptions for various tasks.
-As depicted in Figure 10, our MemVR excels in accurately identifying question-relevant details within images.
-In contrast, as discussed in Appendix C, Qwen-VL-Chat exhibits occasional difficulties in producing detailed image descriptions when applied to VCD.
-These shortcomings become evident in scenarios requiring a nuanced interpretation of image content.
-This suggests that MemVR demonstrates superior adaptability across different architectures, enabling more reliable long-text generation.
+For MME subset evaluations (covering both object-level and attribute-level hallucinations), the results in Figure 2 and Table 8 indicate that MemVR uniformly improves in handling object-level and attribute-level hallucinations, as well as commonsense reasoning.
+The Existence, Count, and Color scores all show significant improvements, where the total score of MME in LLaVA and Qwen-VL increases by 32 and 36, respectively.
+Importantly, although those comparison methods alleviate some hallucinations, they deteriorate generalizability, i.e., they lead to performance degradation on the general benchmarks.
+These results indicate that, compared to CD-based methods, MemVR excels at hallucination mitigation and delivers strong performance on general-purpose benchmarks.
+More experimental results are provided in Appendix B.
+Figure 6: (Left) Comparisons of MemVR with regular LLaVA-1.5, Qwen-VL, and GLM-4V on LLaVA-Bench (in-the-wild).
+(Right) The comparison between MemVR and VCD on LLaVA-Bench shows that MemVR improves significantly while VCD decreases.
+Table 6: LLaVA-Bench (In-the-Wild) evaluation results.
+Methods Convs $\uparrow$ Detail $\uparrow$ Complex $\uparrow$ All $\uparrow$ Average $\uparrow$
+LLaVA-1.5 58.8  $\uparrow$0.0 52.1  $\uparrow$0.0 74.6  $\uparrow$0.0 63.4  $\uparrow$0.0 64.8  $\uparrow$0.0
++ OPERA 59.5  $\uparrow$0.7 49.6  $\downarrow$2.5 78.6  $\uparrow$4.0 59.8  $\downarrow$3.6 64.3  $\downarrow$0.5
++ ICD 40.3  $\downarrow$18. 42.2  $\downarrow$9.9 60.3  $\downarrow$14. 49.8  $\downarrow$13. 56.9  $\downarrow$7.9
++ VCD 57.8  $\downarrow$1.0 50.8  $\downarrow$1.3 77.9  $\uparrow$3.3 59.1  $\downarrow$4.3 63.2  $\downarrow$1.6
++ MemVR 63.8  $\uparrow$5.0 52.6  $\uparrow$0.5 77.9  $\uparrow$3.3 64.0  $\uparrow$0.6 65.2  $\uparrow$0.4
+Table 7: MM-Vet evaluation results, where G denotes language Generation, K: Knowledge S: Spatial awareness, R: Recognition.
+Methods R $\uparrow$ OCR_S $\uparrow$ OCR_K_R $\uparrow$ OCR_G_S $\uparrow$ Total $\uparrow$
+LLaVA-1.5 67.6  $\uparrow$0.0 17.7  $\uparrow$0.0 21.2  $\uparrow$0.0 10.0  $\uparrow$0.0 31.1  $\uparrow$0.0
++ OPERA 61.9  $\downarrow$5.7 21.5  $\uparrow$3.8 11.2  $\downarrow$10. 30.0  $\uparrow$20. 32.0  $\uparrow$0.9
++ ICD 59.5  $\downarrow$8.1 17.7  $\uparrow$0.0 8.8  $\downarrow$12.4 40.0  $\uparrow$30. 25.9  $\downarrow$5.2
++ VCD 62.2  $\downarrow$5.4 15.8  $\downarrow$1.9 17.5  $\downarrow$3.7 60.0  $\uparrow$50. 30.2  $\downarrow$1.1
++ MemVR 70.3  $\uparrow$2.7 23.8  $\uparrow$6.1 21.2  $\uparrow$0.0 30.0  $\uparrow$20. 32.4  $\uparrow$1.3
+Table 8: Performance evaluation on MME Hallucination subset, MM-Vet, Vizwiz, and MMBench. The best results are in green.
+Evaluation Methods MME-Hall Object-Level Attribute-Level MM-Vet Vizwiz MMBench
+Total $\uparrow$ Existence $\uparrow$ Count $\uparrow$ Position $\uparrow$ Color $\uparrow$ Accuracy $\uparrow$ Accuracy $\uparrow$ Accuracy $\uparrow$
+LLaVA-1.5 Regular 643.3  $\uparrow$0.0 190.0  $\uparrow$0.0 155.0  $\uparrow$0.0 128.3  $\uparrow$0.0 170.0  $\uparrow$0.0 31.10  $\uparrow$0.0 50.00  $\uparrow$0.0 62.80  $\uparrow$0.0
+OPERA 610.0  $\downarrow$33. 195.0  $\uparrow$5.0 128.3  $\downarrow$26. 121.7  $\downarrow$6.6 165.0  $\downarrow$5.0 32.00  $\uparrow$0.9 50.76  $\uparrow$0.8 62.80  $\uparrow$0.0
+ICD 583.3  $\downarrow$60. 185.0  $\downarrow$5.0 130.0  $\downarrow$25. 121.7  $\downarrow$6.6 146.7  $\downarrow$23. 25.90  $\downarrow$5.2 37.62  $\downarrow$12. 39.78  $\downarrow$23.
+VCD 648.3  $\uparrow$5.0 190.0  $\uparrow$0.0 155.0  $\uparrow$0.0 133.3  $\uparrow$5.0 170.0  $\uparrow$0.0 30.20  $\downarrow$0.9 44.90  $\downarrow$5.1 54.21  $\downarrow$8.6
+MemVR (ours) 648.3  $\uparrow$5.0 190.0  $\uparrow$0.0 155.0  $\uparrow$0.0 133.3  $\uparrow$5.0 170.0  $\uparrow$0.0 32.40  $\uparrow$1.3 51.50  $\uparrow$1.5 63.75  $\uparrow$0.9
+Qwen-VL Regular 618.3  $\uparrow$0.0 175.0  $\uparrow$0.0 140.0  $\uparrow$0.0 123.3  $\uparrow$0.0 180.0  $\uparrow$0.0 49.00  $\uparrow$0.0 66.05  $\uparrow$0.0 56.53  $\uparrow$0.0
+OPERA  -   $\uparrow$0.0  -   $\uparrow$0.0  -   $\uparrow$0.0  -   $\uparrow$0.0  -   $\uparrow$0.0  -   $\uparrow$0.0  -   $\uparrow$0.0  -   $\uparrow$0.0
+ICD 616.7  $\downarrow$1.7 170.0  $\downarrow$5.0 138.3  $\downarrow$1.7 148.3  $\uparrow$25. 160.0  $\downarrow$20. 31.70  $\downarrow$17. 29.37  $\downarrow$36. 13.32  $\downarrow$43.
+VCD 648.3  $\uparrow$30. 175.0  $\uparrow$0.0 130.0  $\downarrow$10. 153.3  $\uparrow$30. 190.0  $\uparrow$10. 34.60  $\downarrow$14. 34.54  $\downarrow$31. 39.18  $\downarrow$17.
+MemVR (ours) 638.3  $\uparrow$20. 185.0  $\uparrow$10. 145.0  $\uparrow$5.0 123.3  $\uparrow$0.0 185.0  $\uparrow$5.0 49.60  $\uparrow$0.6 66.36  $\uparrow$0.3 56.44  $\downarrow$0.1
+5.4 Ablation Studies
+To validate the effectiveness across the factors we introduced in our MemVR, we conducted in-depth ablation experiments as detailed in Figure 7 and Table 10.
+All ablation experiments are conducted on LLaVA-1.5-7B, and the performance is assessed across benchmarks.
+Impact of Threshold and Injection Ratio.
+Fig.
+7 shows the performance under different injection ratios $\alpha$ and thresholds $\gamma$, we find that setting $\gamma$ between 0.6 and 0.95 can improve performance, with the optimal threshold around 0.75.
+Understandably, a high threshold makes VR hard to trigger, while a low threshold would trigger VR at an earlier layer.
+For injection rate $\alpha$, we find a positive effect of between 5% and 35%, and a negative effect above 35%, which suggests that there is an upper limit to the supplementation of visual memory.
+To drop the hyperparameter $\alpha$ and derive the dynamic injection ratio, we calculate $\alpha$ by $\alpha=2(u-\gamma)$, where $u$ denotes uncertain score and this variant named MemVR, and the results are present in Table 9.
+Figure 7: (Left) Results under different thresholds $\gamma$. (Right) Results under different injection ratios $\alpha$, both on MME benchmark.
+Table 9: MemVR performance on different benchmarks.
+Methods MME (Overall) POPE (COCO) Hallusion -Bench LLaVA -Bench POPE (GQA)
+LLaVA-1.5 1864.68 83.49 41.45 64.80 83.73
++ MemVR 1896.72 88.50 42.34 65.17 89.60
++ MemVR 1894.14 88.40 42.27 65.87 89.57
+Figure 8: Visualisation of uncertainty across layers without and with MemVR. MemVR effectively reduces uncertainty after 8th layer.
+Figure 9: A case study in long text generation. Compared with VCD, MemVR could effectively mitigate hallucinations.
+Impact of VR Strategy.
+MemVR triggers VR dynamically using layer entropy.
+Tab.
+10 shows that dynamic VR outperforms layer-fixed strategy, achieving the highest total score on MME benchmark, where Static-# denotes executing VR on the # layer and Static-$\phi$ means running VR on the specific layer with best performance.
+This indicates that dynamic VR leveraging layer entropy offers a more effective mechanism, adapting better to different scenarios and achieving optimal performance compared to static VR.
+Table 10: MemVR performance with different VR strategies.
+MME Static-7 Static-15 Static-23 Static-$\phi$ Dynamic
+Cognition 347.1 352.14 357.9 362.9 383.9
+Perception 1500.5 1529.1 1500.3 1526.4 1512.8
+Total score 1847.6 1881.2 1858.1 1889.2 1896.7
+5.5 Inference Latency
+MemVR operates dynamically based on uncertainty.
+It uses VR when layer uncertainty exceeds threshold $\gamma$; if uncertainty stays low across all layers, indicating high model confidence, it’s not triggered.
+This mechanism enables efficient inference without extra computation.
+Different from CD-based and Att-intervention paradigms, which need multi-round inferences or have rollback-induced exponential overheads, our proposed look-twice mechanism only requires one regular inference.
+The comparisons on latency, throughput, time cost, and memory are shown in Table 1.
+5.6 Cases Study
+MemVR reduces uncertainty during inference.
+As shown in Figure 8, MemVR effectively reduces uncertainty after the 8th layer triggered VR, further confirming our analysis about the pattern of hallucinations in Section §3.3.
+Long-text capability.
+Beyond single-word QA benchmark evaluation, we explore models’ capacity for comprehensive long-text generation in various tasks.
+As shown in Figure 9, MemVR can accurately identify image details relevant to questions.
+In contrast, as detailed in Appendix B, Qwen-VL-Chat struggles to generate detailed image descriptions in VCD, especially when nuanced image interpretation is needed.
+This indicates MemVR has better cross-architecture adaptability and more reliable long-text generation ability.
+Analysis of failure cases.
+We have collected failure cases from the MME benchmark, in the ’Celebrity,’ ’Scene,’ and ’Landmark’ sub-tasks, where MemVR underperforms compared to the default model.
+We categorize MemVR’s failure cases into two types.
+Type 1 occurs when the default model is right but MemVR is wrong, due to over-disturbing the default model’s reasoning.
+In these instances, the original visual features are sufficient for reasoning, and the reinjected tokens inadvertently disrupt this process, leading to errors.
+We’re exploring ways to reduce such disturbances.
+For failure type 2, it happens when both are wrong, caused by image complexity or MLLM’s original knowledge gaps.
 5.7 Limitations and Further Discussions
-While MemVR demonstrates significant promise in improving the performance in MLLMs, it is not without limitations.
-One major challenge lies in the variability of MLLM architectures.
-Different MLLMs employ varying FNNs, activation functions, and knowledge systems, making it difficult to identify the optimal hyperparameters—such as injection ratios $\alpha$, and premature layer for each specific model.
-This requires considerable effort in model-specific tuning, which may limit the scalability of MemVR without further automation or standardization in hyperparameter selection.
-Additionally, although our research focused on MLLMs with image inputs, MemVR is theoretically applicable to LLMs with different modalities.
-However, we have yet to explore its performance on inputs such as voice, point clouds, 3D meshes, or video frames.
+While MemVR shows substantial potential, it is not devoid of limitations.
+A primary challenge resides in the complex task of identifying the optimal hyperparameters.
+These hyperparameters include the injection ratio of visual information, and the strategy for selecting the triggered layers.
+This represents the focal point of our subsequent research efforts, as we work to perfect MemVR.
+Additionally, although our research focused on MLLMs with visual inputs, theoretically, MemVR can be expanded to more modalities, e.g., listen-twice for audio, scan-twice for spatial perception, check-twice for fMRI, etc.
 This opens up an exciting avenue for future work, where we plan to extend MemVR’s framework to these diverse input formats and assess its efficacy across a broader range of tasks.
-Furthermore, we will explore integrating MemVR into the training procedures of MLLMs, rather than limiting its application to inference, to evaluate whether this could lead to even greater improvements in model performance and generalization.
 6 Conclusion
-This paper proposes a novel training-free paradigm to mitigate hallucination, named MemVR.
-In contrast to previous CD-based methods, which primarily focus on eliminating biases of language priors, MemVR seeks to replenish question-relevant visual clues towards more evidential answers, which signifies the other side of the coin.
-Our experiments, conducted on seven benchmarks, demonstrate the effectiveness of MemVR in mitigating hallucination and improving general performance.
-    A: Ethic Considerations
-  • Appendix
-    B: Theoretical Analysis of MemVR in MLLMs
-  • Appendix
-    C: Additional Experiments, Results, and Discussions
-Appendix A Ethic Considerations
-We list some key ethical considerations of our method:
-Bias and fairness.
-By injecting visual features from CLIP into the Feed-Forward Network (FFN) layers of LLMs, there’s a potential for inherited biases from the original models.
-CLIP, like many pre-trained models, may contain biases in how it represents certain objects, scenes, or demographics.
-These biases could propagate, affecting performance on different types of data (e.g., gender, ethnicity, cultural contexts).
-It’s important to evaluate how MemVR performs across diverse datasets and ensure that it doesn’t reinforce harmful stereotypes or disproportionately fail for certain groups.
-Misuse of Enhanced Models.
-As MemVR aims to improve the long-text generation and overall performance of VLMs, enhanced models could be misused to generate deceptive or misleading content, such as deepfakes or disinformation.
-It’s important to consider whether there are safeguards in place to prevent malicious use of these improved models in scenarios like automated misinformation campaigns or unethical surveillance.
-Data Privacy.
-If the benchmarks used for evaluating MemVR include datasets with personally identifiable information or sensitive content, care should be taken to ensure data privacy.
-Models should be evaluated on publicly available, anonymized, or ethically sourced datasets to avoid violating privacy laws or ethical norms.
-Appendix B Theoretical Analysis of MemVR in MLLMs
-In Multimodal Large Language Models (MLLMs), hallucinations often arise due to insufficient alignment between visual inputs and the model’s internal representations.
+This paper proposes a novel decoding paradigm to mitigate hallucination, named MemVR, and present static and dynamic VR strategies that shift hidden states of the intermediate layer in MLLM for self-enhancement, rather than modulating logits directly with CD manner, thus avoiding multi-round decoding.
+Our experiments, conducted on eight benchmarks, demonstrate the effectiveness of MemVR in mitigating hallucination and improving general performance.
+Importantly, MemVR is a plug-and-play and task-agnostic method compatible with any MLLM, without extra fine-tuning, emphasizing its widespread applicability.
+Acknowledgement
+This work is sponsored by CAAI-Ant Group Research Fund; Guangdong Provincial Department of Education Project (Grant No.2024KQNCX028); Scientific Research Projects for the Higher-educational Institutions (Grant No.2024312096), Education Bureau of Guangzhou Municipality; Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2025A03J3957), Education Bureau of Guangzhou Municipality.
+Impact Statement
+This paper proposes MemVR to alleviate hallucinations in MLLMs.
+The work has potential wider implications.
+Ethically, it may inherit biases from pre-trained models such as CLIP, risking unfair representation.
+There’s also concern that it could be misused to generate disinformation.
+In terms of societal impact, MemVR can improve system reliability in safety-critical areas such as healthcare.
+Appendix A Theoretical Analysis of MemVR
+In Multimodal Large Language Models, hallucinations often arise due to insufficient alignment between visual inputs and the model’s internal representations.
 This paper provides a rigorous theoretical analysis demonstrating that re-injecting visual features into the intermediate layers of MLLMs mitigates hallucinations and enhances representation capability.
@@
 for $l>1$. This inequality indicates that in deeper layers, $H_{vq}^{(l)}$ contains less vision-related information.
-Theorem B.1.
+Theorem A.1.
 Assume that each Transformer layer acts as a deterministic or stochastic mapping with the Markov property.
@@
 Visual Retracing in MLLMs. We reinject vision tokens $Z_{v}$ on $l$-th layer ($ahead\_layer\leq l<L$):
-$\hat{H}_{vq}^{(l)}=\operatorname{FFN}_{\circlearrowright}^{(l)}(H_{vq}^{(l)}\propto Z_{v}).$
+$\hat{H}_{vq}^{(l)}=\operatorname{FFN}^{(l)}(H_{vq}^{(l)}\propto Z_{v}).$
 MemVR ensures that after the $l$-th layer, $\hat{H}_{vq}^{(l)}$ again contains question-aligned visual information.
-Theorem B.2.
+Theorem A.2.
 Let ${H}_{vq}$ be the hidden states of FFN and $\hat{H}_{vq}$ be after reinjection of visual evidence $Z_{v}$.
@@
 ∎
-Theorem B.3.
+Theorem A.3.
 Let $Y$ be the target output dependent on hidden states.
@@
 $H(Y\mid H_{vq}^{(l)})=H(Y)-I(Y;H_{vq}^{(l)}).$
-From Theorem B.2: $\hat{H}_{vq}^{(l)}$ contains more information about $Z_{v}$, i.e., $I(\hat{H}_{vq}^{(l)};Z_{v})\geq I(H_{vq}^{(l)};Z_{v})$.
+From Theorem A.2: $\hat{H}_{vq}^{(l)}$ contains more information about $Z_{v}$, i.e., $I(\hat{H}_{vq}^{(l)};Z_{v})\geq I(H_{vq}^{(l)};Z_{v})$.
 There is $I(Y;\hat{H}_{vq}^{(l)})\propto I(\hat{H}_{vq}^{(l)};Z_{v})$, thus we have $I(\hat{H}_{vq}^{(l)};Y)\geq I(H_{vq}^{(l)};Y)$.
@@
 ∎
-Theorem B.4.
+Theorem A.4.
 Within the Information Bottleneck (IB) framework, reinjecting $Z_{v}$ at layer $m$ optimizes the objective function:
@@
 We denote the maximum possible increase as $\Delta I_{X}^{\max}$.
-For $\Delta I_{Y}$, From Theorem B.2, $\Delta I_{Y}\geq 0$, and suppose we can establish a minimum increase $\Delta I_{Y}^{\min}>0$.
+For $\Delta I_{Y}$, From Theorem A.2, $\Delta I_{Y}\geq 0$, and suppose we can establish a minimum increase $\Delta I_{Y}^{\min}>0$.
 $I(H_{vq}^{(l)};Z_{v}\mid X_{vq})$ represents supplement information about $Z_{v}$ in $H_{vq}^{(l)}$ that is not already explained by $X_{vq}$, and we denote this maximum as $I_{\max}^{Z\mid X}$.
@@
 These theoretical findings provide strong support for the practice of MemVR in MLLMs to improve their performance and reliability.
-Appendix C Additional Experiments, Results, and Discussions
-C.1 Benchmarks
-In this appendix, we provide additional details into the benchmarks referenced in the main paper. To evaluate hallucinations, we employ the following five benchmarks:
-CHAIR Rohrbach et al.
-(2018) evaluates how well the generated captions align with the content of the given image.
-CHAIR consists of two versions: CHAIR_S, which measures the inaccuracies at the sentence level, and CHAIR_I, which evaluates at the object level within the sentence by comparing the number of false objects to the total number of objects.
-For evaluation, we use the val2014 split of the MSCOCO Lin et al.
-(2014) dataset, which includes annotations for 80 object categories.
+Appendix B Additional Experiments, Results, and Discussions
+B.1 Benchmarks and Metrics
+Datasets.
+To rigorously assess the effectiveness of our proposed method, we conduct a comprehensive set of experiments across two benchmarks specifically designed to evaluate hallucination mitigation and five general-purpose benchmarks to gauge the general performance: ① Hallucination benchmarks: Polling-based Object Probing Evaluation (POPE) (Li et al., 2023b), and Caption Hallucination Assessment with Image Relevance (CHAIR) (Rohrbach et al., 2018), HallusionBench (Guan et al., 2024); ② General-purpose benchmarks: VizWiz-VQA (Gurari et al., 2018), MLLM Comprehensive Evaluation (MME) (Fu et al., 2023), Multimodal Benchmark (MMBench) (Liu et al., 2023b), Multimodal Veterinarian (MM-Vet) (Yu et al., 2024b), LLaVA-Bench (in-the-wild) (Liu et al., 2024b).
+In this appendix, we provide additional details on the benchmarks referenced in the main paper. To evaluate hallucinations, we employ the following five benchmarks:
+CHAIR (Rohrbach et al., 2018) evaluates how well the generated captions align with the content of the given image.
+CHAIR consists of two versions: CHAIRs, which measures the inaccuracies at the sentence level, and CHAIRi, which evaluates at the object level within the sentence by comparing the number of false objects to the total number of objects.
+For evaluation, we use the val2014 split of the MSCOCO (Lin et al., 2014) dataset, which includes annotations for 80 object categories.
 We randomly select 500 images from the entire dataset and used the prompt “Please describe this image in detail.” for the MLLM.
-Polling based Object Probing Evaluation (POPE) Li et al.
-(2023c) is a VQA-based metric proposed to assess hallucinations in MLLMs.
+The CHAIR metric includes per-instance evaluation (CHAIRi) and per-sentence evaluation (CHAIRs), defined as follows:
+$\text{CHAIR}_{i}=\frac{|\{\text{hallucinated objects}\}|}{|\{\text{all objectsmentioned}\}|}$
+$\text{CHAIR}_{s}=\frac{|\{\text{sentences with hallucinated object}\}|}{|\{\text{ all sentences}\}|}$
+Polling based Object Probing Evaluation (POPE) (Li et al., 2023b) is a VQA-based metric proposed to assess hallucinations in MLLMs.
 This metric evaluates the MLLM’s response to the prompt “Is [object] is in this image?” To emphasize that this is a binary VQA task, we appended the prompt with “Please answer yes or no.” To select objects referenced in the question prompt, we followed three different sampling options: random, popular, and adversarial.
 We evaluated performance across all sampling options.
-MLLM Evaluation (MME) Fu et al.
-(2023) evaluates the capabilities of MLLMs, dividing the evaluation into two major categories: perception and cognition.
+MLLM Evaluation (MME) (Fu et al., 2023) evaluates the capabilities of MLLMs, dividing the evaluation into two major categories: perception and cognition.
 The perception category includes fine-grained tasks such as existence, count, location, rough color, poster, celebrity, scene, landmark, artwork identification, and OCR.
@@
 All questions in this benchmark are structured to be answered with a simple yes or no.
-Using the LLaVA-Bench Liu et al.
-(2024), we further demonstrated how well our proposed method maintains the language model performance.
+Using the LLaVA-Bench (Liu et al., 2024b), we further demonstrated how well our proposed method maintains the language model performance.
 This benchmark involves posing various situational questions, such as dialogue, detailed descriptions, and complex reasoning, to randomly selected images from the MSCOCO val2014 dataset.
@@
 The generated answers are evaluated by comparing them to the responses of a text-only GPT-4 model.
-C.2 Reproducibility
+Candidate Layers.
+In dynamic premature layer selection, we partition transformer layers into buckets and select one bucket as the candidate layer set.
+For 32-layer LLaVA-1.5-7B, we use two buckets:[0,15),[15,31).
+This design limits the hyperparameter search space to only 2-4 validation runs.
+For efficiency, we use a validation set (MME) to select the best bucket.
+B.2 Backbones and Baselines
+To evaluate our method, we utilize well-known MLLMs: LLaVA-1.5 (Liu et al., 2024b), Qwen-VL (Bai et al., 2023), and GLM4V (Wang et al., 2023), LLaVA-Next (Li et al., 2024).
+Further, We compare our methods with classic training-tree SOTA methods designed to mitigate object hallucination, including visual contrastive decoding SOTA VCD (Leng et al., 2024), OPERA (Huang et al., 2024a) based on overconfidence penalty and hindsight allocation.
+As Dola (Chuang et al., 2023) is layer-wise contrastive decoding for LLMs and performs poorly in MLLMs, it will not be shown in the experiment.
+Experimental results are obtained and benchmarked using unified implementation.
+Greedy search is used as the default decoding strategy in MemVR for all benchmarks.
+For benchmarks, annotation questions are adapted to MLLM templates.
+For POPE, COCO, A-OKVQA, and GQA are used, while MMBench_DEV_EN is used for MMBench.
+MM-Vet is assessed using MM-Vet Online Evaluator, and gpt4-1106-preview is used for LLaVA-Bench.
+CHAIR uses images from COCO Val2014 with the query ”Please describe this image in detail”.
+In MemVR, do_sample=False, temperature=0, threshold=0.75, beam=1.
+All settings of the compared method follow the default configurations from the original papers.
+B.3 Reproducibility
 Implementation details.
 We employed greedy search as the default decoding strategy across all benchmark evaluations.
-For the hallucination benchmarks (POPE and CHAIR) and general-purpose benchmarks (MME, VizWiz-VQA, MMBench, MM-Vet, and LLaVA-Bench (in-the-wild)), questions from the annotation files were used as prompts, formatted to fit the chat templates of each respective MLLM.
+For the hallucination benchmarks (POPE, CHAIR and HallusionBench) and general-purpose benchmarks (MME, VizWiz-VQA, MMBench, MM-Vet, and LLaVA-Bench (in-the-wild)), questions from the annotation files were used as prompts, formatted to fit the chat templates of each respective MLLM.
 Specifically, we utilized the COCO, A-OKVQA, and GQA datasets for POPE evaluation, and MMBench_DEV_EN for MMBench.
-In the MM-Vet evaluation, we used an online evaluator powered by OpenAI GPT-4 to assess generated results, while for LLaVA-Bench (in-the-wild), we employed OpenAI’s model gpt4-1106-preview via API.
+In the MM-Vet evaluation, we used an online evaluator powered by OpenAI GPT-4 to assess generated results, while for LLaVA-Bench (in-the-wild) and HallusionBench, we employed OpenAI’s model gpt4-1106-preview and GPT-4o-mini respectively via API.
 For CHAIR, a randomly sampled image set from the COCO Val2014 dataset was used across all three models, with the prompt ”Please describe this image in detail.” We sampled three different sets of images using different random seeds and evaluated performance by calculating the mean and standard deviation of the results.
@@
 The computational bottleneck was not the numerical accuracy values but the collection of potential hallucinatory factors for analytical purposes, including logits and attention values for each head and layer.
-C.3 Case Study
+B.4 Case Study
 This case study aims to evaluate and present various benchmark cases across multiple domains systematically.
-Figure 11: A case study comparing the levels of hallucination among various baselines
-Figure 12: A case study comparing the levels of hallucination among various baselines
-Figure 13: A case study comparing the levels of hallucination among various baselines
-Figure 14: A case study comparing the levels of hallucination among various baselines
-Figure 15: A case study comparing the levels of hallucination among various baselines
-Figure 16: A case study comparing the levels of hallucination among various baselines
-Figure 17: A case study comparing the levels of hallucination among various baselines
-Figure 18: A case study comparing the levels of hallucination among various baselines
-Figure 19: A case study comparing the levels of hallucination among various baselines
-C.4 Additional Experiments and Results
+Figure 10: A case study comparing the levels of hallucination among various baselines.
+Figure 11: A case study comparing the levels of hallucination among various baselines.
+Figure 12: A case study comparing the levels of hallucination among various baselines.
+Figure 13: A case study comparing the levels of hallucination among various baselines.
+Figure 14: A case study comparing the levels of hallucination among various baselines.
+Figure 15: A case study comparing the levels of hallucination among various baselines.
+Figure 16: A case study comparing the levels of hallucination among various baselines.
+Figure 17: A case study comparing the levels of hallucination among various baselines.
+Figure 18: A case study comparing the levels of hallucination among various baselines.
+B.5 Additional Experiments and Results
+Method    MME
+Overall Perception Cognition
+LLaVA1.5-7B 1864.68  $\uparrow$0.0 1508.97  $\uparrow$0.0 355.71  $\uparrow$0.0
++ VCD (Leng et al., 2024) 1872.87  $\uparrow$8.2 1515.01  $\uparrow$6.0 357.86  $\uparrow$2.2
++ OPERA (Huang et al., 2024a) 1784.34  $\downarrow$80.3 1473.62  $\downarrow$35.3 310.71  $\downarrow$45.0
++ ICD (Wang et al., 2024) 1594.77  $\downarrow$269.9 1306.91  $\downarrow$202.1 287.86  $\downarrow$67.9
++ MemVR (Ours) 1896.72  $\uparrow$32.0 1512.80  $\uparrow$3.8 383.92  $\uparrow$28.2
+Qwen-VL-Chat 1784.93  $\uparrow$0.0 1442.79  $\uparrow$0.0 342.14  $\uparrow$0.0
++ VCD (Leng et al., 2024) 1721.06  $\downarrow$63.9 1403.17  $\downarrow$39.6 317.89  $\downarrow$24.3
++ OPERA (Huang et al., 2024a) - - -
++ ICD (Wang et al., 2024) 1833.75  $\uparrow$48.8 1472.06  $\uparrow$29.3 361.69  $\uparrow$19.6
++ MemVR (Ours) 1820.95  $\uparrow$36.0 1473.45  $\uparrow$30.7 347.50  $\uparrow$5.4
+GLM-4V-9B 2160.54  $\uparrow$0.0 1680.89  $\uparrow$0.0 479.64  $\uparrow$0.0
++ VCD (Leng et al., 2024) 2105.53  $\downarrow$55.0 1624.10  $\downarrow$56.8 481.43  $\uparrow$1.8
++ OPERA (Huang et al., 2024a) - - -
++ ICD (Wang et al., 2024) 2074.30  $\downarrow$86.2 1566.09  $\downarrow$114.8 508.21  $\uparrow$28.6
++ MemVR (Ours) 2170.17  $\uparrow$9.6 1683.03  $\uparrow$2.1 487.14  $\uparrow$7.5
+Table A1: Results on the MME dataset. Best-performing method per model size is in bold. Arrows indicate improvement ( $\uparrow$) or degradation ( $\downarrow$) vs. the baseline model.
+Method     LLaVABench (in-the-wild)
+Average All_1 All_2 All_3
+LLaVA1.5-7B 64.80  $\uparrow$0.0 63.40  $\uparrow$0.0 80.20  $\uparrow$0.0 50.80  $\uparrow$0.0
++ VCD (Leng et al., 2024) 63.20  $\downarrow$1.6 59.10  $\downarrow$4.3 82.00  $\uparrow$1.8 48.50  $\downarrow$2.3
++ OPERA (Huang et al., 2024a) 64.30  $\downarrow$0.5 59.80  $\downarrow$3.6 83.30  $\uparrow$3.1 49.80  $\downarrow$1.0
++ ICD (Wang et al., 2024) 56.90  $\downarrow$7.9 49.80  $\downarrow$13.6 80.70  $\uparrow$0.5 40.20  $\downarrow$10.6
++ MemVR (Ours) 65.17  $\uparrow$0.4 64.00  $\uparrow$0.6 80.20  $\uparrow$0.0 51.30  $\uparrow$0.5
+Qwen-VL-Chat 68.50  $\uparrow$0.0 70.40  $\uparrow$0.0 79.30  $\uparrow$0.0 55.80  $\uparrow$0.0
++ VCD (Leng et al., 2024) 53.77  $\downarrow$14.7 41.00  $\downarrow$29.4 85.30  $\uparrow$6.0 35.00  $\downarrow$20.8
++ OPERA (Huang et al., 2024a) - - - -
++ ICD (Wang et al., 2024) 56.60  $\downarrow$11.9 45.30  $\downarrow$25.1 85.70  $\uparrow$6.4 38.80  $\downarrow$17.0
++ MemVR (Ours) 69.50  $\uparrow$1.0 69.50  $\downarrow$0.9 82.00  $\uparrow$2.7 57.00  $\uparrow$1.2
+GLM-4V-9B 75.30  $\uparrow$0.0 88.40  $\uparrow$0.0 73.00  $\uparrow$0.0 64.50  $\uparrow$0.0
++ VCD (Leng et al., 2024) 74.23  $\downarrow$1.1 86.70  $\downarrow$1.7 72.80  $\downarrow$0.2 63.20  $\downarrow$1.3
++ OPERA (Huang et al., 2024a) - - - -
++ ICD (Wang et al., 2024) 70.97  $\downarrow$4.3 82.10  $\downarrow$6.3 71.80  $\downarrow$1.2 59.00  $\downarrow$5.5
++ MemVR (Ours) 76.73  $\uparrow$1.4 88.90  $\uparrow$0.5 74.80  $\uparrow$1.8 66.50  $\uparrow$2.0
+Table A2: Results on LLaVABench (in-the-wild) dataset.
+Best-performing method per model size and dataset is highlighted in bold; arrows indicate improvement or degradation over the baseline, where higher values indicate better performance.
+Method Total
+LLaVA1.5-7B 31.1  $\uparrow$0.0
++ VCD (Leng et al., 2024) 30.2  $\downarrow$0.9
++ OPERA (Huang et al., 2024a) 32.0  $\uparrow$0.9
++ ICD (Wang et al., 2024) 25.9  $\downarrow$5.2
++ MemVR (Ours) 32.4  $\uparrow$1.3
+Qwen-VL-Chat 49.0  $\uparrow$0.0
++ VCD (Leng et al., 2024) 34.6  $\downarrow$14.4
++ OPERA (Huang et al., 2024a) -
++ ICD (Wang et al., 2024) 31.7  $\downarrow$17.3
++ MemVR (Ours) 49.6  $\uparrow$0.6
+GLM-4V-9B 63.4  $\uparrow$0.0
++ VCD (Leng et al., 2024) 59.4  $\downarrow$4.0
++ OPERA (Huang et al., 2024a) -
++ ICD (Wang et al., 2024) 57.7  $\downarrow$5.7
++ MemVR (Ours) 65.0  $\uparrow$1.6
+Table A3: Results on MM-Vet dataset. Best-performing method per model size is highlighted in bold. Arrows indicate improvement ( $\uparrow$) or degradation ( $\downarrow$) over the baseline.
+Method Accuracy
+LLaVA1.5-7B 50.00  $\uparrow$0.0
++ VCD (Leng et al., 2024) 44.90  $\downarrow$5.1
++ OPERA (Huang et al., 2024a) 50.76  $\uparrow$0.8
++ ICD (Wang et al., 2024) 37.62  $\downarrow$12.4
++ MemVR (Ours) 51.50  $\uparrow$1.5
+Qwen-VL-Chat 66.05  $\uparrow$0.0
++ VCD (Leng et al., 2024) 34.54  $\downarrow$31.5
++ OPERA (Huang et al., 2024a) -
++ ICD (Wang et al., 2024) 29.37  $\downarrow$36.7
++ MemVR (Ours) 66.36  $\uparrow$0.3
+GLM-4V-9B 57.39  $\uparrow$0.0
++ VCD (Leng et al., 2024) 48.04  $\downarrow$9.4
++ OPERA (Huang et al., 2024a) -
++ ICD (Wang et al., 2024) 50.01  $\downarrow$7.4
++ MemVR (Ours) 58.00  $\uparrow$0.6
+Table A4: Results on the Vizwiz dataset.
+Best-performing method per model size is highlighted in bold; arrows indicate improvement ( $\uparrow$) or degradation ( $\downarrow$) relative to the baseline.
+Method     CHAIRS
+Cs Ci Recall Len
+LLaVA1.5-7B 47.60  $\uparrow$0.0 13.30  $\uparrow$0.0 80.60  $\uparrow$0.0 99.70  $\uparrow$0.0
++ VCD (Leng et al., 2024) 55.00  $\uparrow$7.4 15.80  $\uparrow$2.5 77.40  $\downarrow$3.2 101.20  $\uparrow$1.5
++ OPERA (Huang et al., 2024a) 47.60  $\uparrow$0.0 13.50  $\uparrow$0.2 79.00  $\downarrow$1.6 93.20  $\downarrow$6.5
++ ICD (Wang et al., 2024) 56.20  $\uparrow$8.6 16.30  $\uparrow$3.0 16.30  $\downarrow$64.3 103.40  $\uparrow$3.7
++ MemVR (Ours) 46.60  $\downarrow$1.0 13.00  $\downarrow$0.3 80.80  $\uparrow$0.2 99.60  $\downarrow$0.1
+Qwen-VL-10B 6.80  $\uparrow$0.0 5.30  $\uparrow$0.0 53.40  $\uparrow$0.0 17.60  $\uparrow$0.0
++ VCD (Leng et al., 2024) 13.00  $\uparrow$6.2 12.30  $\uparrow$7.0 47.90  $\downarrow$5.5 115.70  $\uparrow$98.1
++ OPERA (Huang et al., 2024a) - - - -
++ ICD (Wang et al., 2024) 18.40  $\uparrow$11.6 14.30  $\uparrow$9.0 37.60  $\downarrow$15.8 48.10  $\uparrow$30.5
++ MemVR (Ours) 4.80  $\downarrow$2.0 3.30  $\downarrow$2.0 52.30  $\downarrow$1.1 15.00  $\downarrow$2.6
+GLM-4V-9B 40.40  $\uparrow$0.0 9.00  $\uparrow$0.0 72.70  $\uparrow$0.0 218.20  $\uparrow$0.0
++ VCD (Leng et al., 2024) 42.20  $\uparrow$1.8 9.60  $\uparrow$0.6 72.80  $\uparrow$0.1 239.80  $\uparrow$21.6
++ OPERA (Huang et al., 2024a) - - - -
++ ICD (Wang et al., 2024) 43.40  $\uparrow$3.0 9.10  $\uparrow$0.1 73.60  $\uparrow$1.9 239.80  $\uparrow$21.6
++ MemVR (Ours) 39.40  $\downarrow$1.0 9.00  $\uparrow$0.0 70.70  $\downarrow$2.0 214.00  $\downarrow$4.2
+Table A5: Results on CHAIRS dataset.
+Best-performing method per model size and dataset is highlighted in bold; arrows indicate improvement or degradation over the baseline, where lower values indicate better performance.
+Method Existence Count Position Color Scene Artwork OCR Numerical Text_trans Code_reason
+LLaVA-Next (Llama3-8B) 195.0 165.0 143.3 185.0 161.6 159.2 118.0 125.0 50.0 77.5
++MemVR 195.0 170.0 143.3 185.0 163.6 161.0 124.0 125.0 52.5 77.5
+LLaVA-Next (Mistral-7B) 190.0 150.0 133.3 190.0 144.2 163.5 113.0 122.5 60.0 67.5
++MemVR 195.0 155.0 133.3 190.0 145.2 165.0 113.8 122.5 60.0 67.5
+LLaVA-Next (Vicuna-1.6-7B) 195.0 135.0 143.3 165.0 162.2 123.2 132.5 42.5 107.5 55.0
++MemVR 195.0 135.0 135.0 170.0 163.0 123.5 140.0 42.5 115.0 57.5
+Table A6: Performance comparison across different LLaVA-Next models with and without MemVR.
+Method     MMBench
+AR CP FP-C FP-S LR RR Overall
+LLaVA1.5-7B 72.86  $\uparrow$0.0 75.68  $\uparrow$0.0 58.04  $\uparrow$0.0 63.48  $\uparrow$0.0 28.81  $\uparrow$0.0 51.30  $\uparrow$0.0 62.80  $\uparrow$0.0
++ VCD (Leng et al., 2024) 60.30 68.58 51.75 53.24 18.64 48.70 54.21
++ OPERA (Huang et al., 2024a) 69.85 75.00 56.64 66.21 28.81 53.04 62.80
++ ICD (Wang et al., 2024) 42.21  $\downarrow$30.7 52.36  $\downarrow$23.3 52.36  $\downarrow$5.7 39.59  $\downarrow$23.9 16.10  $\downarrow$12.7 32.17  $\downarrow$19.1 39.78  $\downarrow$23.0
++ MemVR (Ours) 71.86  $\uparrow$1.2 76.69  $\uparrow$1.0 57.34  $\downarrow$0.7 64.16  $\uparrow$0.9 31.36  $\uparrow$2.5 56.52  $\uparrow$5.2 63.75  $\uparrow$0.9
+Qwen-VL-10B 60.30  $\uparrow$0.0 71.28  $\uparrow$0.0 45.45  $\uparrow$0.0 62.80  $\uparrow$0.0 28.81  $\uparrow$0.0 38.26  $\uparrow$0.0 56.53  $\uparrow$0.0
++ VCD (Leng et al., 2024) 34.67 52.36 20.28 55.63 11.86 22.61 39.18
++ OPERA (Huang et al., 2024a) - - - - - - -
++ ICD (Wang et al., 2024) 12.56  $\downarrow$47.7 17.57  $\downarrow$53.7 2.10  $\downarrow$43.4 22.53  $\downarrow$40.3 2.54  $\downarrow$26.3 5.22  $\downarrow$33.0 13.32  $\downarrow$43.2
++ MemVR (Ours) 61.31  $\uparrow$1.0 71.28  $\uparrow$0.0 44.06  $\downarrow$1.4 62.80  $\uparrow$0.0 27.97  $\downarrow$0.8 38.26  $\uparrow$0.0 56.44  $\downarrow$0.1
+GLM-4V-9B 88.44  $\uparrow$0.0 86.49  $\uparrow$0.0 69.93  $\uparrow$0.0 85.67  $\uparrow$0.0 66.10  $\uparrow$0.0 85.22  $\uparrow$0.0 82.39  $\uparrow$0.0
++ VCD (Leng et al., 2024) 86.43 85.47 68.53 84.64 61.86 81.74 80.58
++ OPERA (Huang et al., 2024a) - - - - - - -
++ ICD (Wang et al., 2024) 83.92  $\downarrow$4.5 84.46  $\downarrow$2.0 62.94  $\downarrow$7.0 81.23  $\downarrow$4.4 60.17  $\downarrow$5.9 80.00  $\downarrow$5.2 78.01  $\downarrow$4.4
++ MemVR (Ours) 88.94  $\uparrow$0.5 86.49  $\uparrow$0.0 70.63  $\uparrow$0.7 86.01  $\uparrow$0.4 66.10  $\uparrow$0.0 85.22  $\uparrow$0.0 82.65  $\uparrow$0.3
+Table A7: Results on the MMBench dataset with newly added ICD rows.
+Best-performing method per model size and dataset is highlighted in bold; arrows indicate improvement ( $\uparrow$) or degradation ( $\downarrow$) over the baseline.
 Strategy 1-Token Len 5-Token Len 10-Token Len 20-Token Len 30-Token Len 50-Token Len 80-Token Len
@@
 Table A8: Time cost for generating tokens. All based on LLaVA1.5-7B
-Method     LLaVABench (in-the-wild)
-Average All_1 All_2 All_3
-LLaVA1.5-7B 64.80  $\uparrow$0.0 63.40  $\uparrow$0.0 80.20  $\uparrow$0.0 50.80  $\uparrow$0.0
-+ VCD (Leng et al., 2024) 63.20  $\downarrow$1.6 59.10  $\downarrow$4.3 82.00  $\uparrow$1.8 48.50  $\downarrow$2.3
-+ OPERA (Huang et al., 2024a) 64.30  $\downarrow$0.5 59.80  $\downarrow$3.6 83.30  $\uparrow$3.1 49.80  $\downarrow$1.0
-+ MemVR (Ours) 65.17  $\uparrow$0.4 64.00  $\uparrow$0.6 80.20  $\uparrow$0.0 51.30  $\uparrow$0.5
-Qwen-VL-Chat 68.50  $\uparrow$0.0 70.40  $\uparrow$0.0 79.30  $\uparrow$0.0 55.80  $\uparrow$0.0
-+ VCD (Leng et al., 2024) 53.77  $\downarrow$14.7 41.00  $\downarrow$29.4 85.30  $\uparrow$6.0 35.00  $\downarrow$20.8
-+ OPERA (Huang et al., 2024a) - - - -
-+ MemVR (Ours) 69.50  $\uparrow$1.0 69.50  $\downarrow$0.9 82.00  $\uparrow$2.7 57.00  $\uparrow$1.2
-GLM-4V-9B 75.30  $\uparrow$0.0 88.40  $\uparrow$0.0 73.00  $\uparrow$0.0 64.50  $\uparrow$0.0
-+ VCD (Leng et al., 2024) 74.23  $\downarrow$1.1 86.70  $\downarrow$1.7 72.80  $\downarrow$0.2 63.20  $\downarrow$1.3
-+ OPERA (Huang et al., 2024a) - - - -
-+ MemVR (Ours) 76.73  $\uparrow$1.4 88.90  $\uparrow$0.5 74.80  $\uparrow$1.8 66.50  $\uparrow$2.0
-Table A2: Results on LLaVABench (in-the-wild) dataset. Best-performing method per model size and dataset is highlighted in bold; arrows indicate improvement or degradation over the baseline.
-Method Total
-LLaVA1.5-7B 31.1  $\uparrow$0.0
-+ VCD (Leng et al., 2024) 30.20  $\downarrow$0.9
-+ OPERA (Huang et al., 2024a) 32  $\uparrow$0.9
-+ MemVR (Ours) 32.4  $\uparrow$1.3
-Qwen-VL-Chat 49.0  $\uparrow$0.0
-+ VCD (Leng et al., 2024) 34.60  $\downarrow$14.4
-+ OPERA (Huang et al., 2024a) -
-+ MemVR (Ours) 49.6  $\uparrow$0.6
-GLM-4V-9B 63.4  $\uparrow$0.0
-+ VCD (Leng et al., 2024) 59.40  $\downarrow$4.0
-+ OPERA (Huang et al., 2024a) -
-+ MemVR (Ours) 65.0  $\uparrow$1.6
-Table A3: Results on MM-Vet dataset.
-Best-performing method per model size and dataset is highlighted in bold; arrows indicate improvement or degradation over the baseline, where lower values indicate better performance.
-Method Accuracy
-LLaVA1.5-7B 50.00  $\uparrow$0.0
-+ VCD (Leng et al., 2024) 44.90  $\downarrow$5.1
-+ OPERA (Huang et al., 2024a) 50.76  $\uparrow$0.8
-+ MemVR (Ours) 51.50  $\uparrow$1.5
-Qwen-VL-Chat 66.05  $\uparrow$0.0
-+ VCD (Leng et al., 2024) 34.54  $\downarrow$31.5
-+ OPERA (Huang et al., 2024a) -
-+ MemVR (Ours) 66.36  $\uparrow$0.3
-GLM-4V-9B 57.39  $\uparrow$0.0
-+ VCD (Leng et al., 2024) 48.04  $\downarrow$9.4
-+ OPERA (Huang et al., 2024a) -
-+ MemVR (Ours) 58.00  $\uparrow$0.6
-Table A4: Results on Vizwiz dataset.
-Best-performing method per model size and dataset is highlighted in bold; arrows indicate improvement or degradation over the baseline, where lower values indicate better performance.
-Method     CHAIRS
-Cs Ci Recall Len
-LLaVA1.5-7B 47.60  $\uparrow$0.0 13.30  $\uparrow$0.0 80.60  $\uparrow$0.0 99.70  $\uparrow$0.0
-+ VCD (Leng et al., 2024) 55.00  $\uparrow$7.4 15.80  $\uparrow$2.5 77.40  $\downarrow$3.2 101.20  $\uparrow$1.5
-+ OPERA (Huang et al., 2024a) 47.60  $\uparrow$0.0 13.50  $\uparrow$0.2 79.00  $\downarrow$1.6 93.20  $\downarrow$6.5
-+ MemVR (Ours) 46.60  $\downarrow$1.0 13.00  $\downarrow$0.3 80.80  $\uparrow$0.2 99.60  $\downarrow$0.1
-GLM-4V-9B 40.40  $\uparrow$0.0 9.00  $\uparrow$0.0 72.70  $\uparrow$0.0 218.20  $\uparrow$0.0
-+ VCD (Leng et al., 2024) 42.20  $\uparrow$1.8 9.60  $\uparrow$0.6 72.80  $\uparrow$0.1 239.80  $\uparrow$21.6
-+ OPERA (Huang et al., 2024a) - - - -
-+ MemVR (Ours) 39.40  $\downarrow$1.0 9.00  $\uparrow$0.0 70.70  $\downarrow$2.0 214.00  $\downarrow$4.2
-Qwen-VL-10B 6.80  $\uparrow$0.0 5.30  $\uparrow$0.0 53.40  $\uparrow$0.0 17.60  $\uparrow$0.0
-+ VCD (Leng et al., 2024) 13.00  $\uparrow$6.2 12.30  $\uparrow$7.0 47.90  $\downarrow$5.5 115.70  $\uparrow$98.1
-+ OPERA (Huang et al., 2024a) - - - -
-+ MemVR (Ours) 4.80  $\downarrow$2.0 3.30  $\downarrow$2.0 52.30  $\downarrow$1.1 15.00  $\downarrow$2.6
-Table A5: Results on CHAIRS dataset.
-Best-performing method per model size and dataset is highlighted in bold; arrows indicate improvement or degradation over the baseline, where lower values indicate better performance.
-Method     MMBench-Dev-EN
-AR CP FP-C FP-S LR RR Overall
-LLaVA1.5-7B 72.86  $\uparrow$0.0 75.68  $\uparrow$0.0 58.04  $\uparrow$0.0 63.48  $\uparrow$0.0 28.81  $\uparrow$0.0 51.30  $\uparrow$0.0 62.80  $\uparrow$0.0
-+ VCD (Leng et al., 2024) 60.30 68.58 51.75 53.24 18.64 48.70 54.21
-+ OPERA (Huang et al., 2024a) 69.85 75.00 56.64 66.21 28.81 53.04 62.80
-+ MemVR (Ours) 71.86  $\uparrow$1.2 76.69  $\uparrow$1.0 57.34  $\downarrow$0.7 64.16  $\uparrow$0.9 31.36  $\uparrow$2.5 56.52  $\uparrow$5.2 63.75  $\uparrow$0.9
-GLM-4V-9B 88.44  $\uparrow$0.0 86.49  $\uparrow$0.0 69.93  $\uparrow$0.0 85.67  $\uparrow$0.0 66.10  $\uparrow$0.0 85.22  $\uparrow$0.0 82.39  $\uparrow$0.0
-+ VCD (Leng et al., 2024) 86.43 85.47 68.53 84.64 61.86 81.74 80.58
-+ OPERA (Huang et al., 2024a) - - - - - - -
-+ MemVR (Ours) 88.94  $\uparrow$0.5 86.49  $\uparrow$0.0 70.63  $\uparrow$0.7 86.01  $\uparrow$0.4 66.10  $\uparrow$0.0 85.22  $\uparrow$0.0 82.65  $\uparrow$0.3
-Qwen-VL-10B 60.30  $\uparrow$0.0 71.28  $\uparrow$0.0 45.45  $\uparrow$0.0 62.80  $\uparrow$0.0 28.81  $\uparrow$0.0 38.26  $\uparrow$0.0 56.53  $\uparrow$0.0
-+ VCD (Leng et al., 2024) 34.67 52.36 20.28 55.63 11.86 22.61 39.18
-+ OPERA (Huang et al., 2024a) - - - - - - -
-+ MemVR (Ours) 61.31  $\uparrow$1.0 71.28  $\uparrow$0.0 44.06  $\downarrow$1.4 62.80  $\uparrow$0.0 27.97  $\downarrow$0.8 38.26  $\uparrow$0.0 56.44  $\downarrow$0.1
-Table A6: Results on MMBench dataset.
-Best-performing method per model size and dataset is highlighted in bold; arrows indicate improvement or degradation over the baseline, where lower values indicate better performance.
-C.5 Supplement Implement Detail
-The code of VCD Leng et al.
-(2024) is also released.
-However, the result of VCD evaluated in our experiments (e.g.
-POPE and MME benchmarks) is lower than the original paper.
+B.6 VCD Implement Details
+The code of VCD (Leng et al., 2024) is also released.
+However, the result of VCD evaluated in our experiments on POPE is lower than the original paper.
 Therefore, we report the results in the original paper.
-Appendix D Examples of capability integrations
-Table A7: Six samples on MM-Vet benchmark requiring different capability integrations.
+Appendix C Examples of capability integrations
 [Uncaptioned image] [Uncaptioned image]
@@
 Required capabilities: Recognition, OCR, Spatial Awareness, Math Required Capabilities: OCR, Knowledge, Spatial Awareness
-D.1 Others
+Table A9: Six samples on MM-Vet benchmark requiring different capability integrations.
+C.1 Others
 Q: How can the model understand information directly from the vision encoder, especially if it has a different vision system?
 To ensure that MemVR is adaptable across diverse vision systems, we conducted experiments on multiple VLM architectures, including LLaVA, which utilizes a Visual-Instructional-Tuning framework with different sizes of ViT-based CLIP models, Qwen-VL-Chat, which employs a Q-Former-like architecture for visual processing, and ChatGLM-4v-9B, which integrates a large pre-trained visual encoder.
-These architectures encompass a broad range of vision models, providing confidence that MemVR is applicable to most VLMs in use today.
-Artifacts and licenses
-We report a list of licenses for all datasets and models used in our experiment in Table A8. We strictly follow all the model licenses and limit the scope of these models to academic research only.
+These architectures encompass a broad range of vision models, providing confidence that MemVR is applicable to most MLLMs in use today.
+Artifacts and licenses We report a list of licenses for all datasets and models used in our experiment in Table A10.
+We strictly follow all the model licenses and limit the scope of these models to academic research only.
 Data Sources URL License